### v2

# Deterministic and Reliable Software-Defined Vehicles: key building blocks, challenges, and vision

Abstract
As vehicle systems become increasingly complex, with more features, services, sensors, actuators, and processing units, it is important to view vehicles not just as modes of transportation moving toward full autonomy, but also as adaptive systems that respond to the needs of their occupants. Vehicular services can be developed to support these adaptations. However, the increasing complexity of vehicular service development, even with current standardizations, best practices and guidelines, are insufficient to tackle the high complexity of development, with expectations of up to 1 (U.S.) billion lines of code for a fully (level 5) autonomous vehicle.
Within this survey, the paradigm of Deterministic Software Defined Vehicles is explored, aiming to enhance the quality and ease of developing automotive services by focusing on service-oriented architectures, virtualization techniques, and the necessary deterministic intra- and inter-vehicular communications. Considering the main open challenges for such verticals, a vision architecture towards improved services development and orchestration is presented, focusing on: a) a deterministic network configurator; b) a data layer configurator; c) a hypervisor configurator; d) the vehicle abstraction layer; and e) a software orchestrator.
keywords:
Software-Defined Vehicles, Connected Vehicles, Vehicular Applications, Vehicular Services, Intelligent Transportation Systems, Deterministic Networks, Time Sensitive Networks, Survey
journal: Computer Communications
[…]
[overall]organization=Department of Electronics, Telecommunications and Informatics, University of Aveiro / Instituto de Telecomunicações Aveiro, addressline=Campus Universitário de Santiago, city=Aveiro, postcode=3810-193, country=Portugal
[…]
[duarte]organization=Instituto de Telecomunicações Aveiro, addressline=Campus Universitário de Santiago, city=Aveiro, postcode=3810-193, country=Portugal
3GPP
Third Generation Partnership Project
5QI
5G QoS Identifier
AADL
Architecture Analysis and Design Language
ABS
Anti-lock Braking System
ACM
Association for Computing Machinery
ADAS
Advanced Driving Assistance System
ANACOM
Autoridade Nacional de Comunicações
AP
Access Point
API
Application Programming Interface
AR
Augmented Reality
ASIC
Application-Specific Integrated Circuit
ASIL
Automotive Safety Integrity Level
ASN.1
Abstract Syntax Notation One interface description language
ATCLL
Aveiro Tech City Living Lab
ATS
Asynchronous Traffic Shaping
AUTOSAR
Automotive Open System Architecture
AVB
Audio Video Bridging
Bosch
Bosch Car Multimédia Portugal
Bosch Ovar
Bosch Security Systems, S.A
BSM
Basic Safety Message
CAM
Cooperative Awareness Message
CAN
Controller Area Network bus
CAN-FD
CAN Flexible Data Rate
CAV
Connected Autonomous Vehicle
CBS
Credit Based Shaper
CCU
Central Control Unit
CNC
Centralized Network Controller
CNI
Container Network Interface
CoAP
Constrained Application Protocol
COVESA
Connected Vehicle Systems Alliance
CPM
Collective Perception Message
CPS
Collective Perception Service
CPU
Central Processing Unit
CSMA-CA
CSMA-Collision Avoidance
CSMA
Carrier-Sense Multiple Access
CUC
Centralized User Configuration
CUDA
Compute Unified Device Architecture
C-V2X
Cellular Vehicle-to-Everything
DC-GBR
Delay Critical GBR
DCU
Data Collection Unit
SDP
Simple Discovery Protocol
DDS
Data Distribution Service
DENM
Decentralized Environmental Notification Message
DetNet
Deterministic Networking
DL
Deep Learning
DoS
Denial of Service
DPDK
Data Plane Development Kit
DRY
Don’t repeat yourself
DSL
Domain-Specific Language
DSP
Digital Signal Processor
DSRC
Dedicated Short Range Communications
eBPF
extended Berkeley Packet Filter
ECU
Electronic Control Unit
E/E
Electrical/Electronic
ELK
Elasticsearch-Logstash-Kibana
ETSI
European Telecommunications Standards Institute
EVA
Emergency Vehicle Alert
eVTOL
electrical Vertical Take Off and Landing
FACE
Future Architecture for Computing Environment
FDMA
Frequency-Division Multiple Access
FPGA
Field-Programmable Gate Array
FRER
Frame Replication and Elimination for Reliability
GBR
Guaranteed Bit Rate
GDA
Geographical Destination Area
GNSS
Global Navigation Satellite System
GPRS
General Packet Radio Service
GPS
Global Positioning System
GPU
Graphical Processing Unit
gRPC
Google RPC
HMD
Head Mounted Display
HPC
High Performance Computer
HTTP
Hypertext Transfer Protocol
HUD
Heads-Up Display
I2X
Infrastructure to Everything
IC
Integrated Circuit
ICA
Intersection Collision Announcement
IEEE
Institute of Electrical and Electronics Engineers
IETF
Internet Engineering Task Force
IMU
Inertial Measurement Unit
I/O
Input/Output
IoT
Internet of Things
IoU
Intersection over Union
IP
Internet Protocol
IPFS
InterPlanetary File System
IT-Aveiro
Institute of Telecommunications of Aveiro
ITS-S
Intelligent Transport System Station
ITS
Intelligent Transport System
JDL
U.S. Joint Directors of Laboratories
JSON
JavaScript Object Notation
KPI
Key Performance Indicator
KVM
Kernel-based Virtual Machine
Lidar
LIght Detection And Ranging
LIN
Local Interconnect Network bus
LoRaWAN
Long Range Wide-Area Network
LPWAN
Low-Power Wide-Area Network
LTE
Long Term Evolution
LTE-A
Long Term Evolution Advanced
M2M
Machine to Machine
MAC
Media Access Control
MANET
Mobile Ad-hoc Network
MCM
Manoeuvre Coordination Message
MEC
Multi-access Edge Computing
MISRA
Motor Industry Software Reliability Association
ML
Machine Learning
MOST
Media Oriented Systems Transport
MPLS
Multiprotocol Label Switching
MQTT
Message Queuing Telemetry Transport
MTBF
Mean Time Between Failures
MTU
Maximum Transmission Unit
NAP
Network Architectures and Protocols
NB-IoT
Narrowband IoT
NETCONF
Network Configuration Protocol
NFC
Near Field Communication
NGSIv2
Next Generation Service Interfaces version 2
NIC
Network Interface Controller/Card
NOMA
Non-orthogonal Multiple Access
NTP
Network Time Protocol
OASIS
Organization for the Advancement of Structured Information Standards
OATD
Open Access Theses and Dissertations
OBD2
OnBoard Diagnostic v2
OBU
On-Board Unit
OEM
Original Equipment Manufacturer
OFDMA
Orthogonal FDMA
OMG
Object Management Group
OOP
Object Oriented Programming
OS
Operating System
OSAL
Operating system abstraction layer
OSI
Open Systems Interconnection
OTA
Over-the-air
P2I
Pedestrian to Infrastructure
P2P
Pedestrian to Pedestrian
P2V
Pedestrian to Vehicle
P2X
Pedestrian to Everything
PCP
Priority Code Point
PER
Packet Error Rate
POSIX
Portable Operating System Interface
PPS
Pulse Per Second
PSM
Personal Safety Message
PTP
Precision Time Protocol
QoS
Quality of Service
Radar
RAdio Detection And Ranging
RAP
Resource Allocation Protocol
RAT
Radio Access Technology
RDMA
Remote Direct Memory Access
REST
Representational State Transfer
RESTCONF
NETCONF through REST
RFC
Request for Comments
ROI
Region of Interest
ROS
Robotic Operating System
RPC
Remote Procedure Call
RSU
Road Side Unit
RTE
Run Time Environment
RTOS
Real-Time Operating System
RTPS
Real-Time Publish/Subscribe Protocol
RTT
Round Trip Time
rTWT
restricted Target Wake Time
SAE
Society of Automotive Engineers
SCS
Stream Classification Service
SDK
Software Development Kit
SDN
Software-Defined Network
SDR
Software-Defined Radio
SDV
Software-Defined Vehicle
SDVC
Software-Defined Vehicular Cloud
SIP
Session Initiation Protocol
SLA
Service Level Agreement
SOA
Service Oriented Architecture
SOAFEE
Scalable Open Architecture for Embedded Edge
SoC
System-on-chip
SOME/IP
Scalable service-Oriented MiddlewarE over IP
SPICE
Automotive Software Process Improvement and Capability Determination
SPoF
Single Point of Failure
SRP
Stream Reservation Protocol
TAS
Time Aware Shaper
TCP
Transmission Control Protocol
TDD
Test Driven Development
TDMA
Time-Division Multiple Access
TIM
Traveler Information Message
TSF
Timing Synchronization Function
TSN
Time Sensitive Network
TTC
Time To Collision
TTEthernet
Time-Triggered Ethernet
TTL
Time To Live
DETI/UA
Department of Electronics, Telecommunications and Informatics, University of Aveiro
UC
Use Case
UDP
User Datagram Protocol
UE
User Equipment
UI
User Interface
UPF
User Plane Function
U.S.
United States of America
USB
Universal Serial Bus
UTP
Unshielded Twisted Pair cable
V2I
Vehicle to Infrastructure
V2P
Vehicle to Pedestrian
V2V
Vehicle to Vehicle
V2X
Vehicle to Everything
VAL
Vehicle Abstraction Layer
VAM
Vulnerable Road User Awareness Message
VANET
Vehicular Ad-hoc Network
VCU
Vehicle Control Unit
VM
Virtual Machine
VPN
Virtual Private Network
VRU
Vulnerable Road User
VXCAN
Virtual CAN tunnel
WAN
Wide Area Network
WAVE
IEEE 802.11p WAVE
WPAN
Wireless Personal Area Network
WGS84
EPSG:4326/World Geodetic System 1984
WSGI
Web Server Gateway Interface
XDP
eXpress Data Path
XML
eXtended Markup Language
YANG
Yet Another Next Generation
YOLO
You Only Look Once real-time object detection system
1 Introduction
[…]
Figure 1: Evolution of complexity in vehicular architectures [4].
This growing code base is becoming unmanageable to vehicle manufacturers and developers of vehicular services within a SDV. As described in fig. 2, automakers will tend to leave unchecked, unverified and poorly tested software go into production vehicles. If the growth trend of the code base continues, the results can be potentially deadly [^3] Unverified logic and deadly effects on Tesla’s autopilot per washingtonpost.com.. The overall intra-vehicular systems, as developed following legacy standards and practices, are deemed too complex and static — and thus costly to maintain, update and keep safe — by vehicle manufacturers, with the reduction of the overall complexity (both in software and hardware terms, in particular wiring harness weight savings) becoming a major goal in order to reduce operational costs and decrease potential security issues from unchecked, poorly tested code [5].
Standardization of best practices and development guidelines for intra vehicular software gained traction during the early 2000s [6]. In addition, programming methodologies and guidelines [^4] Such as the ASIL within ISO 26262 or MISRA-C, a set of guidelines for safety-related systems, such as automotive vehicles. Details at synopsys.com, misra.org.uk and synopsys.com. are critical and cannot be disregarded. However, they are no longer enough to tackle such complexity of development.
The context of deterministic pieces of software that carry out something within the vehicle or its environment can provide a solution to this problem. Vehicular services, since they may be critical, shall be deterministic in a broad sense in order to ensure that the result is exactly what was intended in the exact moment it was intended. In this article, we propose that the quality and easiness of the development of services for automotive can be improved using the paradigm of Deterministic Software Defined Vehicles. The main contributions of this survey are as follows:
  1. 1.
    Demonstrate the suitability and relevance of the software defined vehicles for a sustainable development of vehicular services.
  2. 2.
    Improve the state-of-the-art on Software Defined Vehicle concept and its main building blocks.
  3. 3.
    Propose an integration of Time-Aware and Deterministic Networks concepts and solutions within a Software Defined Vehicle.
  4. 4.
    Definition of a set of potential use cases that aim to frame the proposed integration not only in the foreseeable future, but in a long-term perspective.
  5. 5.
    Survey of hardware and software alternatives for validation of the proposed integration.
The remaining of this document is organized into the following sections. Section 2 introduces the background concepts, defining vehicular services and their characteristics, and the concept of SDV and additional information on the industrial landscape. It also shows how automakers or suppliers are facing the SDV paradigm. Next, section 3 presents the main building blocks for a SDV based vehicle (and vehicular services) development, and section 4 discusses on the main challenges. Section 5 follows with the main use cases and requirements that can be drawn from the use cases. Section 6 considers those requirements for the proposal of an architecture and of an evaluation process to validate the work. This considers the required hardware and software for both simulation and real-world testing. Finally, conclusions are provided in section 7.
Figure 2: Average complexity of individual software projects grown by 300 percent over the past decade, creating an unsustainable gap between software complexity and the productivity of software developers [3, 7].
2 Software Defined Vehicles
At the core of SDV there is the concept that features within a vehicle are mostly defined through software. This follows a trend where vehicles, that started with little to no features, then evolved to features defined by mechanical components and by the 1980s started to define the features through software. Figure 3 summarizes this evolution.
Figure 3: Evolution of vehicular features over time, including levels of SDV. Partially based on [8, 9].
The SDV paradigm main ideas are not new. Prior to the advent of SDVs, the automotive industry focus on software was mostly restricted to connectivity between vehicles and the user’s smartphones to provide basic insights into the state of the vehicle [10]. By 1998, [11] considered intelligent vehicles as essential to autonomous transportation, designing a proposal of a vehicle capable of communicating with a centralized management control system. Also in 1998, [12] proposes that smart vehicles are a junction of communication architectures, system architecture, sensors, tasks, kinematics and man-machine interface. The authors propose a communication architecture in a producer-distributor-consumer pattern with buses such as RS232 and Controller Area Network bus (CAN), where time coherence is necessary. By 2001, a reference model architecture for intelligent vehicles [13] considered the vehicle as having intelligent control systems with each layer with one or more computational nodes interconnected and connected to sensors.
Automotive vehicles need to evolve to a new concept of SDV since in-vehicle services must be dynamic, and should be able to adapt over the vehicle life-cycle. At the same time, in-vehicle services need to interact with the vehicle’s systems, sensors, and the surrounding infrastructure within a well-defined time-frame. Therefore, systems must also be interconnected with a high-performance, high-precision, reliable set of networks and protocols with time guarantees.
SDV aims to improve cars along their life-cycle [14], while potentially mitigating the increased unmanaged complexity of modern vehicular software, hardware, and communications. The paradigm states that vehicles manage their operations, add functionality, and enable new features primarily or entirely through software [^5] More details at blackberry.qnx.com.. Since vehicles evolve over their lifetime, with more features, solutions that may require remote control or remote updates, main features being delivered by software adds flexibility, reducing the difficulty of adding new features and allowing new business models to be considered by automakers during the lifetime of the vehicle.
The shift towards SDV means that new features can be pushed to existing and new vehicles, which are now more defined through software than through hardware or mechanical systems as in the past [3]. Not only this adds flexibility to vehicle manufacturers, but also creates new business models where owners can pay for features after buying the vehicle and during its lifetime. This follows a trend where vehicles are now seen by potential owners in a subscription payment model, instead of the one-time payments or leasing/pay-per-use models (fig. 4).
Figure 4: Preferred payments scheme to purchase/activate connectivity services, % of respondents (n = 1649) showcasing a shift in preferred ownership payment models [15].
Such support not just improves the development of vehicular services, but also helps towards a decrease in hardware complexity (contributing towards a much needed decrease in the number of ECUs, cables and the cost and weight savings which follow), with more uniform software code-base (and therefore reduction of the costs and complexity of the development) and greater adaptability (essential specially if the future of the car ownership is car sharing [16]). This adaptability also supports a greater capability to process higher volumes of data while fulfilling the overall performance requirements — essential in Connected Autonomous Vehicles.
2.1 Industrial landscape
Vehicular manufacturers are aware of the necessary changes in thinking. However, the transition from the previous mentality of closed source, customized solutions for each model of vehicle towards a more extendable, interoperable, architecture is far from being over. The Renault, Volvo/Polestar and Volkswagen groups seems to be ahead in terms of a more interoperable architecture based on ECUs interconnected through Ethernet that run Android and well-known versions of chipsets. It is however not possible to ignore efforts from other vehicle manufacturers that may be developing equivalent solutions but have not, at the time of this survey, announced their efforts. The opacity of the vehicular industry, while expected within a competitive market, it remains and will remain a threat to any related work, including this work. To mitigate this threat, research of gray literature is essential, targeting not only automakers but their suppliers and open source community initiatives.
Figure 5 summarizes the efforts of automakers on adoption the main aspects related to SDV, with table 1 providing further detail per associated areas.
Y5.0e0]ptPrimitivevehiclesystemY5.0e0]ptExplosionofvehiclesystemsY9.0e0]ptAUTOSARClassicY5.0e0]ptAUTOSARAdaptiveY2.0e0]ptCARIADBMWY3.0e0]ptRenaultVolvoY6.0e0]ptOpportunity?Y8.0e0]ptTeslaMore featuresMore versatileMore standardMore open
Figure 5: Positioning of main automakers and frameworks over the overall industrial landscape.
Table 1: Overview of the manufacturers effort on SDV.
[t] Entity State of integration Ethernet TSN Zonal ECUs SOA SDV Open Standard AUTOSAR Classic $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ AUTOSAR Adaptive $\times$ $\times$ ✓- ✓- $\times$ BMW $\times$ Renaulta ✓- ✓- Volvo ✓-c $\times$ Volkswagen groupb ✓? ✓? ✓? ✓? $\times$ $\times$ NVIDIA $\times$ $\times$ $\times$ $\times$ Tesla ✓- Eclipse $\times$
Existing automakers were not agile enough within the development of their vehicular services [17], which lead to the rise of new automakers, such as Tesla, in very software-intensive tasks such as autonomous driving. The Tesla company is in fact probably the most well-known example of application of the SDV paradigm principles. Within its vehicles, the ECUs — such as the Autopilot Control Unit, the Media Control Unit, and several Body Control Units — are interconnected through Ethernet, and are in a lower number than the usual within vehicles up to that point. This zonal architecture supports Advanced Driving Assistance System (ADAS) systems for level 2+ autonomous driving, such as the Full Self-Driving (Supervised) that is then sold to end-users in a subscription manner [^6] Details at tesla.com/support/full-self-driving-subscriptions..
As a standardized ”standardized software framework and open E/E system architecture”, the Automotive Open System Architecture (AUTOSAR)777autosar.org is a de facto standard, with the support from vehicle manufacturers and vendors. It is a highly complex ecosystem, with more than 100 modules and its documentation reaching more than 21 000 pages [6]. The standard has two different types of platforms as depicted in fig. 6: the Classic, for traditional hard real-time systems based on micro-controllers with a static configuration, and set of pre-compiled monolithic binaries; and also a newer Adaptive platform with greater flexibility and adaptability. The Classic platform is aimed at safety critical systems with hard real-time requirements with a stated linkage between services and CPU cores. Development through the original Classic Platform is however, very hard-coded to the vehicle and its Electrical/Electronic (E/E) architecture [18]. The Adaptive is capable of considering heterogeneous high performance ECUs and departs from legacy buses [19] such as CAN towards a Service Oriented Architecture (SOA) communication architecture (see  section 3.2) with Ethernet based on Scalable service-Oriented MiddlewarE over IP (SOME/IP) as bus protocol [20] and is overall more dynamic but less time predictable [21, 22], essential within critical vehicular systems.
Figure 6: AUTOSAR Classic VS Adaptive.
After the landscape changing forced by the United States through Tesla, European and Asiatic automakers such as BMW [23]; the Renault group through the Renault Software Factory [^8] More details on www.reuters.com.; Volvo; Kia and Hyundai [^9] More details on www.carscoops.com.; and the Volkswagen group through the CARIAD [^10] Available at cariad.technology., supported by companies such as Vector [24] and VXworks, are shifting towards a software-first development of the vehicles, based on novel hardware and communication paradigms.
A major player within the European market, the Volkswagen group take on SDV is based on creating and controlling the whole platform — cloud (Volkswagen Automotive Cloud VW.AC), hardware and software (VW. Operating System (OS)) [25], affecting interoperability with other automakers and hardware providers. While in initial stages, it highlights a possible tendency towards a close proprietary SDV paradigm. However, it was not possible to obtain documentation on APIs or Software Development Kits that should be used to developed vehicular services using the CARIAD ecosystem. A notable exception is the Roobadge [^11] Product description at https://www.volkswagen.com.au/en/ roobadge.html., a solution to avoid collisions with kangaroos — very common in Australia — by adding, through a front badge or license plate holder sensors to detect the kangaroos. This solution not only showcases a real example of the potential of SDV, since new features are added to an existing vehicle through a new deployed service, but is also compatible with other brand’s vehicles, showcasing a real SDV example that considers brand interoperability (even if the actual implementation is closed).
Still within Europe, the Renault group considers a backbone of ECUs interconnected with Ethernet, with gateways to bridge the Ethernet connected ECUs with legacy sensors or actuators interconnected by CAN [14, 26]. These ECUs run Android and are based on Qualcomm chipsets, supporting an interoperable, open SDV architecture [^12] More details at www.qualcomm.com.. Volvo and Polestar have a similar approach, with zone controllers (more in section 3.1) — high performance controllers in specific zones of the vehicle — interconnected and connected to a central master Vehicle Control Unit (VCU) with an Ethernet backbone [22]. The Zone Controllers can also act as a bridge to legacy buses, like CAN and Local Interconnect Network bus (LIN). Volvo within their SDV efforts also published OpenAPI specifications for their CAVs [27].
Other European players include Mercedes-Benz, BMW and Stellantis. Mercedes-Benz, after a transition to a zonal E/E architecture (with separations between infotainment, body and comfort, ADAS and interior domains), is working on the MB.OS, a proprietary chip-to-cloud platform capable of ensuring decoupling software from hardware. This solution is already allowing Over-the-air (OTA) updates to significantly change the MB User Experience (MBUX), with future focus on standardization efforts [28]. The BMW group also aims to go from signal oriented to unified vehicle wide layered SOA architecture with providers and consumers [23, 29]. This also requires latencies, configurations and bandwidth options encoded in the system. The Stellantis group aims at supporting sustainability and user experience use cases through service oriented solutions. SOA can provide the necessary scalability previous service development paradigms could not — namely 1 application per ECU — while allowing a firm contract between services [30], as described in section 3.2.
Within the United States of America, General Motors is working towards their SDV efforts through the uServices solution, a set of APIs for developers and Ultifi, a end-to-end software platform that “will enable OTA updates and in-car subscription services” [31].
When on the tier-2+ suppliers for automakers, it is clear they are also working towards providing automakers with new tools. Continental, with the Amazon Internet of Things (IoT) for Automotive [^13] More details at catalog.workshops.aws/awsiotforautomotive/. support, is also working towards a “a modular multi-tenant hardware and software framework that connects the vehicle to the cloud” [^14] More details at aws.amazon.com.. Nvidia is entering the market with a clear AI first approach — the Nvidia DRIVE OS, which includes a embedded Real-Time Operating System (RTOS), a hypervisor, Compute Unified Device Architecture (CUDA) libraries and TensorRT, and features such as secure boot, security services, firewall, and OTA updates [^15] Developer documentation at developer.nvidia.com/drive. Vector [^16] At vector.com within its ecosystem propose tools to support development of vehicles with ECUs interconnected by Ethernet with TSN and of services using AUTOSAR within SOA architectures. MotionWise [32] proposes a middleware for safe orchestration of vehicle software across ECUs for automated driving platforms, with support for AUTOSAR classic and adaptive, Robotic Operating System (ROS), and Data Distribution Service (DDS) support. Hardware companies such as Renesas propose the AosCloud [^17] More details at renesas.com., a solution to support Original Equipment Manufacturers and service providers from seamless software installation to operation in the vehicle, and implement various new services and functional enhancements simply.
Regarding the open-source community, alternatives such as the Eclipse Foundation, within its SDV working group, is also working within incubating initiatives to define: a) orchestrators and virtualization managers (Ankaios, Velocitas); b) vehicle abstraction models (Chariott); c) a development platform (Kuksa); and d) an abstraction layer for communication between nodes (eCAL) and operating system (MoEc.OS) with support for DDS as well as simulators such as SUMO and CARLA (see section 6.2). While in an initial development phase, these projects can be a good future basis for anyone entering the SDV. The Scalable Open Architecture for Embedded Edge (SOAFEE)181818soafee.io, more details at arm.com. presents a cloud-native alternative with an architecture enhanced for mixed-criticality automotive applications with corresponding open-source reference implementations to enable commercial and non-commercial offerings. Autoware, an open-source project for autonomous driving191919autoware.org, through its Open AD Kit, the first SOAFEE blueprint, enables cloud-native DevOps of autonomous driving solutions for SDV [^20] Developer documentation at autowarefoundation.github.io..
These individual initiatives are being target towards a cooperation. By 2024, at CES 2024, AUTOSAR, Connected Vehicle Systems Alliance (COVESA), Eclipse SDV, and SOAFEE announced a “collaboration of collaborations” to “align efforts in the SDV ecosystem and create a joint SDV vision clear and unified definition of what constitutes an SDV, which points towards a brighter future in SDV with greater interoperability [33]. Innovation projects such as the HAL4SDV212121hal4sdv.eu also aim to contribute to a more holistic solution by including in its goals a ”standardized interfaces to sensors, actors, compute resources and persistent storage”.
Figure 7: Overview of the topics involved in modern vehicular development to be referred, lightly based on [34].
3 Main building blocks for Software Defined Vehicles
Developing a vehicle within the SDV paradigm ( a ) requires considerations on several domains within the context provided by fig. 7, including:
  • 1.
    Physical disposition of the vehicular ECUs and specifications of each ECU ( c section 3.1): this domain explores how ECUs are physically disposed throughout the vehicle, namely considering a zonal Electrical/Electronic architecture.
  • 2.
    Service Oriented Architectures ( b section 3.2): this domain explores how vehicular services are developed and communicate between one another and abstractions for the vehicular service development and for transferring data between services.
  • 3.
    Virtualization techniques and operating systems ( d , e sections 3.4 and 3.3): this domain considers how vehicular services are deployed within each ECU, including the software stack within each ECU and virtualization techniques.
  • 4.
    Intra and inter vehicular communication architectures, protocols and physical layers ( f sections 3.5 and 3.6 this domain considers how vehicular ECUs communicate with each other and with ECUs on other vehicles (intra-vehicular and inter-vehicular communications, respectively).
  • 5.
    Deterministic vehicular communications ( g section 3.7): this domain considers how to ensure vehicular communications always happen and within the necessary reliability and timing requirements.
The next subsections further detail each one of these building blocks, including their state of the art and associated main challenges.
3.1 Physical E/E architectures
Figure 8: Legacy distributed, domain, and zonal E/E architectures within a vehicle.
To support SDVs, the Electrical/Electronic and computational architectures within vehicles must also evolve from the fully distributed vision, with ECUs scattered throughout the vehicle towards a more centralized solution, with zone-based architectures, i.e. ECUs positioned in specific parts of the vehicle responsible for certain types of actions [35]. Figure 8 provides an overview of this evolution.
During the 1980s (fig. 8a)), new features were added at the cost of adding new specialized ECUs , close to the sensors and actuators and connected through CAN or other equivalent bus. With the explosion of features in vehicles, especially luxury cars, during the 1990s, the number of ECUs increased significantly, making vehicles more complex, and thus increasing the amount of systems that need to be choreographed, developed, tested and that have to communicate, as well as the possibilities of something going wrong [36, 37].
By the 2000s, the high complexity of intra vehicular systems became subject to de facto standards such as AUTOSAR [36, 37] and buses such as FlexRay aimed at mitigating some of the shortcomings of CAN. Since the ECUs were scattered throughout the vehicle, with little to no centralization, these legacy architectures were distributed, allowing a better decoupling, separation of concerns and higher tolerance to equipment failure. However, challenges such as the scalability of the deploying services across multiple ECUs (considering hardware variations and the absense of virtualization), communication performance, and the high costs of the associated complex, bulky, wiring harnesses have driven a trend towards a more centralized architecture, consolidating processing into fewer computing units.
A centralized architecture can reduce costs and complexity by decreasing the number of ECUs, increasing software reusability, and enabling greater flexibility.The centralized E/E architectures supporting SDVs follow the principle of aggregation of ECUs into a small subset of more powerful High Performance Computers. Each HPC can potentially contain multi core Central Processing Units, Graphical Processing Units, Digital Signal Processors, and potentially Field-Programmable Gate Arrays [36, 37, 38], as well as high performance network capabilities (e.g. Ethernet). Some ECUs can eventually be dedicated to computing intensive tasks, including Machine Learning (ML) and Deep Learning (DL) workloads supporting ADAS systems [39]. This follows previous trends within other industries, in particular the aviation industry [19].
Centralization within the intra-vehicular systems can be achieved at a domain, cross domain or zone/vehicle centralized way [19]. By 2010s, domain and cross domain architectures (fig. 8b)) were a first step towards mitigating those issues through the aggregation of functions per domain. In a domain-oriented solution, the grouping of ECUs into HPCs is done by vehicle function since each function requires heavy interaction between ECUs (specially in complex systems such as ADAS). In a cross-domain solution, functions of several domains are consolidated within more powerful ECUs. This minimizes, without removing, the high bandwidth required for communication between ECUs. However, the integration of multiple domains required by systems such as ADAS still require coordination between different domains [36, 37].
Zone-oriented architectures (fig. 8 (b)) aimed to solve the issue of coordination between domains. Each zone of the vehicle collects and forwards information to a much more powerful central ECU, or even offloads to the cloud. Bosch reported on achieving a 15–20% reduction in wiring harness weight when using a zone-oriented architecture instead of a domain-oriented architecture, showcasing the potential of zonal architectures versus the previous approaches.
Future approaches should also consider other vehicles and infrastructure (e.g. smart cities) as an integral part of the SDV environment services can consider for deployment.The next phases of automotive software development should also have a deep integration with cloud and follow the best practices of software development (such as CI/CD) [40] and should consider deterministic networking concepts and standards, as discussed in section 3.7.
3.2 Service-oriented architectures
As vehicular E/E architectures become more complex, standards and well defined patterns are required in order to model both the ECUs hardware and the software. Service Oriented Architectures are an example of such architectural pattern, where each component provide and consume services. Figure 9 presents an overview of the SOA paradigm. Services in this context are a discrete unit of functionality, from directly interacting with sensors and actuators, up to more complex interactions between simpler services. A service contains the mechanism to enable access to one or more capabilities, where the access is provided using a prescribed interface and is exercised consistent with constraints and policies as specified by the service description [41].
Figure 9: SOA in vehicles as perceived by the Renault Group automaker (adapted). Services (in red) are the basic unit of vehicular applications and communicate with one another via one or more message bus [14].
When implemented on a vehicle, on a more reduced set of (zonal) ECUs, the SOA paradigm brings benefits such as the decoupling of hardware and software, increased modularity and separation of concerns, and contribute towards the overall decrease in hardware complexity. [42]. These advantages are, however, at the cost of increased complexity in maintaining contracts between services and ensuring availability and discovery of the services [42]. This is particularly complex within a vehicular context, with a SOA supporting at most 60 to 80 services [23], not addressing situations where there are hundreds or thousands of micro-services.
Within SOA based architectures, data layers (message bus in fig. 9 can abstract the physical medium of services by providing a unified way of exchanging and accessing data between services on different components/ECUs [43]. Information from vehicle sensors can then be fed from/to services leveraging the ECUs and their communication paradigm through data layer solutions.
Three main paradigms can be considered for implementing a data layer [44]: a) a fully periodic publisher-subscriber based or b) a fully event triggered client-server based. Examples for the publish/subscribe paradigms include the DDS [^22] At dds-foundation.org. middleware, SOME/IP, or Message Queuing Telemetry Transport (MQTT[22]. Publish/subscribe paradigms can respond to a data centric paradigm, when the data sending is event triggered, unlike solutions such as Hypertext Transfer Protocol (HTTP) or Constrained Application Protocol (CoAP) based on the request/response paradigm [24]. Abstraction layers such as uprotocol [^23] At projects.eclipse.org/projects/automotive.uprotocol. can be considered to decouple vehicular services from the chosen data layer.
Table 2 provides a summary on some of the main possible solutions for SOA data middleware. While all these solutions can ensure delivery of data, there is a tendency to consider DDS and Zenoh the best performant solutions since they provide better flexibility and performance. According to [45], which compared performance of MQTT, DDS, and Zenoh, the latter shows (for multi-machine scenarios, more realistic in zonal vehicular architectures) a double increase on the throughput of DDS and 50 times of MQTT, while DDS usually provides the best latency and better flexibility with more Quality of Service (QoS) levels.
Table 2: Qualitative comparison of IoT standards (partially based on [46, 47, 48]).
[t] Standard MQTT DDS SOME/IP Zenoh HTTP Transport UDP (in MQTT-SN), TCP UDP, TCP IP UDP, TCP, QUIC TCP, QUIC Paradigm Pub / Sub Pub / Sub Pub / Sub Pub / Sub Data storage Request / Reply Node Discovery No Yes (SDP and alternatives) Yes Yes No Fault Tolerance Broker is SPoF Yes, fully distributed by default Yes Yes, can be fully distributed Client and server are SPoF Data Schema None Required Required Only encoding is required Optionalc QoS Policies At delivery level At multiple entitiesa N/A At 2 levelsb N/A
  • a
    For delivery, presentation, transport, availability, cache, resource management.
  • b
    For reliability and fragmentation level.
  • c
    Based for example on the Content-Type.
3.3 Vehicular operating systems
ECUs within a vehicle run services that require deterministic scheduling from the OS. Failure to have a deterministic real-time scheduling can lead to services being dispatched too soon or too late to the OS, with dramatic effect on the services behavior. Therefore, Real-Time Operating System are a need within vehicular computing domains to ensure services are dispatched in a determined specific order and time-frame. While many OSs can support real-time constraints — including any generic Linux kernel based OS with PREEMPT_RT [^24] Available at wiki.linuxfoundation.org/realtime. patches [49], some solutions are heavily automotive focused, such as Android Automotive, Nvidia Drive OS and QNX.
Android Automotive [^25] Not to be confused with Android Auto, a platform to use a smartphone as a infotainment alternative. and its extensions such as SnapOS [^26] Available at snappautomotive.io. provide an OS and a platform that is based on Android — and can therefore leverage the existing application ecosystem — while supporting specific characteristics of vehicles, including the support for the creation of a vehicle abstraction layer required to abstract the vehicle’s specific/proprietary components, UI support for setting vehicle’s settings and OTA updates.
Nvidia Drive OS [^27] Dev documentation at developer.nvidia.com/drive. is the “reference operating system and associated software stack designed specifically for developing and deploying autonomous vehicle applications on DRIVE AGX-based hardware”. It is in fact a custom RTOS with a Nvidia Hypervisor and necessary CUDA and TensorRT libraries to fully utilize the hardware (graphical) potential.
Yocto [^28] Available at yoctoproject.org. based OSs are also a possibility. Yocto supports the creation of custom OS based on Linux, allowing adding custom layers to embedded systems. Yocto based Wind River Linux [^29] At windriver.com. is an example of such family of OS, providing support for multiple board support packages in both x86 and Arm hardware. Automotive Grade Linux [^30] At docs.automotivelinux.org, github.com/nxp-auto-linux. is another example of a Yocto based OS and a industry supported initiative to develop software for the “infotainment, instrument cluster, Heads-Up Display (HUD), telematics, connected car, ADAS, functional safety, and autonomous driving” [50].
Alternative OSs include the VxWorks [^31] Available at windriver.com. [51], a RTOS for mission critical embedded systems. While usually considered in space scenarios, automotive are also supported for example through the integration with AUTOSAR Adaptive Platform. The QNX [^32] Available at blackberry.qnx.com. is another alternative as a micro-kernel RTOS with extensive foothold in the automotive sector installed in more than 215 million vehicles. Another open source initiative, albeit in initial stages (launched by January 2024), is the Eclipse ThreadX [^33] Previously Azure RTOS, now threadx.io, with dev documentation at eclipse.org., an open-source, vendor neutral RTOS.
These operating systems, being mostly RTOS or RTOS-like, can be evaluated through the worst-case execution time for each task, modularity, response time or jitter [52]. At the time of writing, no comparison between some or all of the aforementioned OSs was found.
Vehicular services should be developed independently of the RTOS and thus an isolation between both is desired. Initiatives such as the Operating system abstraction layer (OSAL) from IBM [^34] Dev documentation at ibm.com. or NASA “isolates embedded application software from a RTOS by providing “a well-defined, generic interface to RTOS services” [^35] More information at opensource.gsfc.nasa.gov. and can, therefore, contribute to the necessary portability of services between OSs. Without such initiatives, the definition of the OS will be mostly related to the associated ecosystem, hindering possiblity interoperability between services developed for each OS.
3.4 Virtualization and orchestration techniques
To ensure optimal usage of ECUs, with concurrent execution of different services with different criticality, partitioning of scheduling of each core per service(s) is required. Virtualization of resources, either memory, buses (e.g. CAN) and CPU allow both isolation of services within virtual machines while enabling migration and reusage of legacy software [19].
Considering the need for flexibility but also real-time constraints, ECUs can leverage virtualization principles to support multiple operating systems, both real-time and general purpose OSs, through virtualization.
Virtualization provides support for running multiple OSs on the same hardware, ensuring the abstraction of hardware and of network topologies [53].One of the key advantages is allowing hardware to be shared and not uniquely assigned to one operating system, while achieving the necessary isolation automakers require between critical and non-critical services. It is an essential concept in the context of vehicular systems not only because of isolation but to support the aggregation of functions in less ECUs (per the zonal architectures) than the number of services and required OSs [54].
Virtualization creates a number of challenges, including:
  • 1.
    Orchestration: how to distribute resources (both location and quantity) of the virtualization continuum across virtual machines.
  • 2.
    Shared I/O: how to distribute input and output devices, such as Universal Serial Bus (USB) or Ethernet buses, across virtual machines.
  • 3.
    Virtualization with real-time performance: since schedulability at hypervisor level is not deterministic [19].
Allowing virtualized services to use and share the same Ethernet port for example requires a scheduler that is capable of communicating with the different virtual machines and the deterministic network configurator. This communication can be achieved through hypercalls, the equivalent of syscalls but to invoke a hypervisor [55]. This may impact, however, the overall frame transmission delays.
When virtualization is considered, having the scheduler of the traffic can be at either application level, virtual machine level or hypervisor level [54]. At applicational level, since there’s one scheduler per application, there are going to be race conditions between each scheduler, with non-deterministic behaviors and therefore providing results with high jitter. At virtual machine level, for example leveraging the Linux Traffic Control subsystem, while the necessary configurations are transparent to the applications, the hypervisor will not be aware of the special traffic. Working at the hypervisor level solves the previous issue through special Hypercalls invoked by the middleware within the Hypervisor level. Hypercalls [^36] Similarly to syscalls concept, hypercalls provide an interface between a guest OS and the hypervisor. may be used within the middleware to support deterministic networkng within virtualized intra vehicular services [55].
Vehicular OSs such as Android Automotive provides full support for virtualization through Kernel-based Virtual Machine (KVM) and virtIO [^37] Developer documentation at wiki.libvirt.org. devices, allowing support at the Virtual Machine (VM) side for sensors, graphics, storage, network (including Wi-Fi through virtio-net and VirtWifi [^38] Developer documentation at android.com.) and with support for communicating with the vehicle through Google RPC (gRPC) [^39] Details at android.com..
Generic Linux-based OSs also support virtualization through containerization. Containerized services have advantages over hypervisor virtualization such as the reduced overhead and reduced storage space usage [39].
Eclipse, within the SDV foundation proposes Eclipse Ankaios [^40] Developer documentation at projects.eclipse.org/projects/automotive.ankaios. to “manage multiple nodes and virtual machines with a single unique API in order to start, stop, configure, and update containers and workloads”. Ankaios is tailored towards automotive applications and developed for automotive ECUs and provides a central place to manage automotive applications, supporting:
Kubernetes [^41] Available at kubernetes.io., as a container orchestration tool mainly designed for datacenters, can also be consider to manage vehicular workloads. The work of [56] proposes vehicles as nodes of a Kubernetes cluster.
Apart from its inherent advantages as container orchestration and scheduler, Kubernetes is very flexible in networking terms, since networking between containers can be changed through Container Network Interfaces [^42] Project available at github.com/containernetworking/cni.. CNIs are a specification for configuring network interfaces in containers. This allows considering deterministic networking concepts (section 3.7) in container scenarios. For example, the work of [57] proposes support for deterministic networking within Kubernetes containers without modifying the application binaries through a side car container that exposes the standard Portable Operating System Interface (POSIX) socket interface to other containers and redirects traffic to deterministic configured network interfaces in the host computer. Not modifying the application binaries is essential to ensure compatibility with existing vehicle services but also to simplify the development process of new services.
3.5 Intra-vehicular communications
Vehicular systems need to communicate with other systems within and outside the vehicle. While most intra-vehicular communication is expected to be through wired physical layers, this does not have to be the norm. Table 3 provides an overview of the main buses that may be used for intra-vehicular communications.
Table 3: Overall comparison of main intra vehicular wired buses [19].
[t] Characteristic / Bus LIN CAN FlexRay Ethernet Media 1 wire Unshielded Twisted Pair cable (UTP) 2/4 wires UTP Topology Bus Bus Star, bus Star, bus Access control Master-slave Priority-based messages TDMA Switched point-to-point link Max. Baud Rate 19.2 Kb/s 1 Mb/s, up 16 Mb/s (XL)a 10 Mb/s 10 Mb/s, 100 Mb/s, 1000+ Mb/s MTU (Bytes) 8 8/64 254 1472 (UDP), 1460 (TCP), 9000 (Jumbo) Max nodes 16 (1 master) 16 (500 Kb/s) 22 (bus), 22/64 (star), 64 (mixed) 256 to theoretically limited Security Modules None SecOC SecOC MACsec, IPsec, TLS/DTLS, SecOC
  • a
    If disregarding recommenced maximum data-rates [2].
[…]
In the past, these limitations could be addressed through static analysis of the overall vehicular systems and ECUs. However, the increasing number of features —particularly data-intensive functionalities related to semi-autonomous driving— has necessitated a significant expansion in the volume and complexity of the wiring harness. This increase involves not only a greater number of wires but also added complexity, leading to a substantial weight of at least 40-50 kg [^45] Source: semiengineering.com..
While legacy buses like CAN or LIN can and will continue to be on edge intra vehicular systems, such as smaller more specific ECUs, sensors and actuators, the long term objective is the edges nodes to be also connected through Ethernet [61]. Within the context of vehicular networks, Ethernet is being more and more considered as the backbone connection between high performance ECUs, with gateways to legacy buses like CAN. As depicted in fig. 10, there are multiple iterations of Ethernet standards, i.e. there’s not a single cable and solution for all scenarios. Automakers shall take into consideration the characteristics of the cables (such as width, weight or resistance to heat and vibrations)— and not only the theoretical bandwidth — in order to choose the best ones to interconnect vehicular ECUs. Several physical layer specifications can be considered, including IEEE 100BASE-T1, 1000BASE-T1, and 1000BASE-RH [62, 63]. Possible Ethernet backbone upgrades to ensure future scalability are also being explored, including 10BASE-T1S or multi-Gb links.
Figure 10: Evolution of Ethernet data rate and standards over time [63]
[…]
Towards the goal of support Ethernet on the vehicles, major automakers and their providers have created the OPEN Alliance (One-Pair Ether-Net) [^46] At opensig.org. initiative. Canonical also defends E/E architectures which combines and consolidates ECUs and interconnects them using an Ethernet network [3]. Such initiatives could also benefit from the resilience proven by industry grade protocols, such as PROFINET [^47] At profibus.com..
Ethernet can be integrated with CAN for example through the CAN-XL protocol. This is done by mapping an Ethernet frame into a CAN-XL frame. Since CAN-XL provides 2048 data bytes, Q-tagged or enveloped Ethernet frames fit into a CAN-XL frame [66]. Examples of gatewaying strategies from CAN to Ethernet has been presented by [67], including FPGA based gateways with suitable latency, routing mechanisms between Ethernet and CAN or FlexRay, with the authors proposing a gatewaying up to the application layer — SOME/IP in this case through a SOME/IP service and a RTOS program that read the CAN messages.
Alternatives to Ethernet include FlexRay and Media Oriented Systems Transport (MOST). Unlike CAN, FlexRay has the unique ability to sync up nodes on a network without an external synchronization clock signal by using two special types of frames [68], being a possible alternative to Ethernet with Deterministic mechanisms (more on section 3.7). The MOST specification defines “the protocol, hardware and software layers necessary to allow for the efficient and low-cost transport of control, real-time and packet data” using “different physical layers (fiber optic, UTP, coax)” and supporting 25, 50 and 150 Mbps data transfer speeds [69], still under the theoretical maximus of Ethernet.
3.6 Inter-vehicular communications
For inter vehicular communications, it is impractical to consider wired physical layer. Within these context, the Vehicle to Everything (V2X) paradigm is of increasing importance considering the advent of both Vehicle to Vehicle (V2V) scenarios and smart cities. The usage of cloud infrastructure is also an integrated part of the SDV paradigm, requiring Vehicle to Infrastructure (V2I) wireless communications. For example, CARIAD, the initiative from the Volkswagen group towards SDVs, launched an uniform application store for all brands within the Volkswagen group [25], that indicates a strong cloud presence within its ecosystem.
Vehicular nodes can use different wireless access technologies to communicate with other road elements, including vehicles, V2V; pedestrians, Vehicle to Pedestrian (V2P); and infrastructure, V2I; creating a Vehicular Ad-hoc Network (VANET), a sub-type of a Mobile Ad-hoc Networks. These technologies can be classified based on their communication range [70], with a summary of the main characteristics of the most used access technologies described in table 4.
Table 4: Example of wireless vehicular communications access technologies and main characteristics.
[…]
Cellular based solutions, either through LTE, Long Term Evolution Advanced (LTE-A) or 5G, can provide connection to other vehicles and infrastructure, with high coverage and medium latency at a high monetary cost [75].
Wi-Fi, including its latest iteration (Wi-Fi 7 / IEEE 802.11be), can also theoretically provide connectivity to and from vehicles, as well as IEEE 802.11bb / Li-Fi. Li-Fi adapts the Media Access Control (MAC) layer of 802.11 while supporting light (visible or infrared) based physical layers in order to overcome interference in dense deployments (which Wi-Fi is particularly susceptible to) and restrictions or inefficiencies of radio frequencies - potentially increasing the maximum. An evolution of the original infrared physical layer proposed in 1999 for IEEE 802.11, it is targeted at the industrial markets [78]. However, both solutions fail to consider the high mobility of vehicular nodes, which means frequent handovers between access points.
Vehicular targeted standards include Cellular Vehicle-to-Everything (C-V2X) and ITS-G5. C-V2X is cellular based and aims at integrating cellular technologies within V2V scenarios. Based on cellular standards it also allows direct device-to-device communication with low latency and high reliability, but in short ranges (±1 km), with low cost associated with the possible necessary hardware changes. C-V2X also support device-to-network, enabling cloud services to be part and parcel of the end-to-end solution. On the other hand, ITS-G5 is based on IEEE 802.11p for physical and MAC layer. It requires existence of road infrastructure — Road Side Units — and that vehicles are equipped with On-Board Units. Problems of this alternative includes handling handovers and scaling to many nodes simultaneously and vulnerability to packet collisions due to concurrent transmissions, and packet collisions due to hidden vehicular nodes. Existing literature shows that C-V2X offers performance advantages over ITS-G5 in terms of its additional link budget, higher resilience to interference, and better non-line-of-sight capabilities [79].
Radio Access Technology (RAT) evolutions, such as the IEEE 802.11bd and New Radio for C-V2X can supplement today’s vehicular networks in supporting higher bandwidth and QoS wireless communications. IEEE 802.11bd are more suitable for supporting more V2X applications efficiently, with better throughput and low transmission latency, by adapting the physical and MAC layer techniques introduced in IEEE 802.11n/ac/ax, while ensuring interoperability and backwards compatibility with 802.11p. [80].
3.7 Deterministic intra and inter-vehicular communications
[…]
Services running through a communication network and with requirements of high performance and high precision should consider not only techniques to decrease the overall end-to-end latency of data from/to services — for example by leveragging the edge computing concept [83, 84] — but also techniques to ensure the latency requirements and real-time hard constraints. The usage of deterministic communications and networks that support such communications are vital when considering vehicular services that are somehow critical. Determinism, per the Merriam-Webster dictionary, “is the quality or state of being determined”, i.e. the quality of being firmly resolved. Within networks this concept can be regarded as the delivery of units of information — either abstracted in frames or packets or other abstraction — in a fully resolved manner in several domains, such as temporal or reliability, in all the steps of the frame/packet from the producer to the consumer(s) of information.
Regarding the temporal determinism, within the exchange of a unit of data between two network entities one may consider the main steps of a packet switching as presented in fig. 11. Each step introduces delay that, if the network is not properly configured, can be unbounded and potentially even random. Even if high bandwidth is a possibility, transmission, and queuing delays still exist, affecting end-to-end delay [79, 81].
Figure 11: Main steps of a packet switching. Partially adapted from [85].
Sets of standards such as TSN [^48] At 1.ieee802.org/tsn. and Deterministic Networks [^49] At datatracker.ietf.org/wg/detnet. can ensure that high reliability and time precision for each step of the packet switching process [86]. Current solutions to tackle these issues within the TSN and Deterministic Networks are further presented in this section. The next subsections present the state-of-the-art on the main initiatives and sets of standards towards determinism in networks in lower Open Systems Interconnection (OSI) layers. These initiatives started for wired mediums such as Ethernet but can also be extended to wireless contexts.
3.7.1 In wired context
Methods towards achieving deterministic communications within a wired intra-vehicular context (such as the connectivity between vehicular ECUs) are being standardised, with two major initiatives to be considered: a) the Institute of Electrical and Electronics Engineers (IEEETSN set of standards, operating at OSI layer 2; and b) the Internet Engineering Task Force (IETF) Deterministic Networking (DetNet) operating at OSI layer 3. Both initiatives shall be seen not as competitors but as complementary initiatives.
The IETF Deterministic Networking set of standards aims at defining deterministic — i.e. with bounded latency, loss and jitter — paths for the packets [87]. Those standards operate at layer 3 and can be supported by a Software-Defined Network (SDN) approach, where one or more controllers are responsible for defining the nodes behavior, to warrant [88]:
  • 1.
    Explicit routes: to ensure deterministic routing. This may be achieved for example through the usage of Multiprotocol Label Switching (MPLS), a routing technique based not on network addresses but labels.
  • 2.
    Resource allocation: to ensure protection against congestion, resources (such as bandwidth) is allocated.
  • 3.
    Service paths: to ensure protection against equipment and medium failures packets are replicated through disjoint paths.
Table 5 provides an overview on the Request for Comments (RFC) standards and documents that define the goals, use cases and building blocks behavior. The DetNet building blocks operate at the Internet Protocol (IP) layer and use lower-layer technologies such as MPLS and TSN. These building blocks can and should be controlled based on observable metrics (hops, their state, and end-to-end latency), and on fault detection/prediction and identification [88].
Table 5: Main DetNet standards, specifications and documents.
Document Goal Importance
RFC 8557 Problem statement Contextualizes the need for DetNet with the need for a deterministic forwarding of packets.
RFC 8578 Use cases Validates self-driving vehicles as examples of use cases that can be addressable by DetNet techniques.
RFC 8655 Architecture Defines the main building blocks of DetNet.
RFC 8938, 8939, 8964, 9023, 9024, 9025, 9037, 9056 Data Plane Defines how building blocks may be implemented and metadata that packets require. Includes information of integration with TSN and MPLS for explicit routing.
RFC 9016 YANG model (draft) [89] Flow information model Modelling of the configuration information for configuration of DetNet building blocks.
RFC 9055 Security Considerations on security, including challenges posed by replication of packets and spoofing.
RFC 9320 Bounded Latency Provides mathematical modelling of the bounded latency values for a DetNet architecture.
At a lower OSI layer, the Time Sensitive Network group presents an effort to solve transmission, and queuing delays issues. Table 6 provides an overview on the IEEE standards and documents that define the goals, use cases and building blocks behavior. TSN standards adds to Ethernet mechanisms to ensure time characteristics, performance, reliability, and security. Mechanisms include time synchronization (e.g. gPTP [82]), replication and elimination for reliability (e.g. Frame Replication and Elimination for Reliability (FRER)) [90], and bounded low latency (e.g. credit-based [91] or time aware shaping algorithms [92]) that can provide the necessary real-time performance and bounding of the latencies [93, 94, 95].
[…]
In conjunction with DetNet, but operating at OSI layer 2 — and originally on top of wired Ethernet networks — the IEEE 802.1 TSN set of standards provide mechanisms for [98]:
  • 1.
    Time synchronization: applications and other network mechanisms require a precise (down to nanosecond accuracy) synchronization between the nodes of the network. Protocols such as Precision Time Protocol (PTP[82] can provide such precision within wired and wireless networks. Synchronization must take into account physical layer latencies and their variations, as these can impact propagation delay measurements and overall synchronization accuracy. IEEE 802.3cx proposes a mechanism to report the physical layer delays in a dynamic way [99].
  • 2.
    Replication and elimination for reliability: improve the reliability by ensuring that frames are replicated and sent through disjoint paths. This replication needs additional mechanisms to ensure repeated frames are ignored. Solutions such as FRER [90] can manage this replication.
  • 3.
    Resource management: Mechanisms to ensure the configuration of TSN talker, listener and of equipment between both and the necessary allocation of resources.
  • 4.
    Bounded latency or bandwidth: ensure the bandwidth or latency and its jitter is deterministic by ensuring transmission happens at a specific time.
Regarding bounded latency, traffic shaping can provide the necessary real-time performance and bounding of the latencies or bandwidth [93, 94, 95], with shapers such as:
  • 1.
    Credit Based Shaper (CBS): based on regulating the transmission rate of frames to smooth out traffic, reduce bursts and ensure bandwidth for certain traffic classes, by delaying the transmission of some frames. When the queue contains a frame to be transmitted, a credit counter is checked and the frame is only transmitted if the counter is non-negative (and if it is, is held back until is non-negative). This counter accumulates credits when the transmission queue is idle, and consumes credits when frames are transmitted, with the rate of accumulation and consumption configurable via parameters (idle slope and send slope) [^50] More details at inet.omnetpp.org. [91].
  • 2.
    Time Aware Shaper (TAS): works by associating the data to traffic classes and the traffic classes to gates that open and close at specific times and for specific intervals of time, ensuring no unwanted competition between traffic classes that should not compete for transmission, therefore eliminating the jitter associated with that competition [92, 100].
  • 3.
    Asynchronous Traffic Shaping (ATS): which is based on the leaky bucket principle, with transmission allowed when credit is greater than the size of the frame [97].
Table 6: Main TSN standards [98] (partially adapted).
Document Goal Importance
802.1Qca Path control and reservation Define mechanisms to allocate resources to transmit frames.
802.1Qbv Time-aware scheduling Defining when and what frames are sent by the emissor.
802.1Qbu, 802.3br Frame preemption Defines the possibility of higher priority frames interrupting lower priority, essential in ensuring precise transmission time for higher priority flows.
802.1Qcc Central configuration model Defines the need for coordination nodes (the CUC and CNC). Updates on the SRP protocol.
802.1Qci Filtering and policing Defines mechanisms to count and filter frames and detect and mitigate disruptive flows of frames.
802.1CB Redundancy through frame replication and elimination Defines mechanisms to replicate frames and manage the repeated frames to increase reliability of the network and avoid equipment failure impacting PER and similar metrics.
1588, 802.1AS Precise ($ns$ level) time synchronization between nodes Defines mechanisms for a nanosecond synchronization between clocks of several network nodes.
P802.1DG (draft) TSN profile for automotive in-vehicle Ethernet communications. Recommendations for TSN implementation within vehicular architectures.
How the network should be configured is also addressed through the IEEE 802.1Qcc standard, which describes three possible models [101]:
  • 1.
    A fully distributed model where all the end stations communicate their requirements — through protocols such as Stream Reservation Protocol (SRP) or Resource Allocation Protocol (RAP).
  • 2.
    A centralized network/distributed user model, where an entity, the CNC, gather the network capabilities.
  • 3.
    A fully centralized model, where the CNC is helped by the CUC to ensure proper network equipment configuration.
These three models are based on the existence of two entities — the Centralized Network Controller (CNC) and the Centralized User Configuration (CUC). The CNC defines the overall network schedule and is the main controller of the network within the network bridges, dispatching configuration with protocols such as Network Configuration Protocol (NETCONF). The CUC mediates the requests between end devices and their applications [102], collects all user requirements and interacts with the CNC to decide the topology and scheduling of the network. Figure 12 provides an overview of these interactions.
Figure 12: Interaction between CUC, CNC and the remaining TSN topology [102].
TSN is not however the only standard at layer 2 for deterministic communications. Time-Triggered Ethernet (TTEthernet), a standard for industrial and avionics applications improved from the Avionics Full Duplex Switched Ethernet, is an alternative[19]. While the concepts of deterministic timing of the transmission of frames is inherently the same, both differ in a number of aspects [103, 104]. TSN is a set of standards while TTEthernet is a specific SAE AS6802 standard. In the former, the schedule is on entire traffic classes, while in latter is on individual frames. In addition, TSN supports lower-priority traffic using non-used bandwidth, while TTEthernet blocks other traffic even when there are no time-scheduled frames to transmit results in a less efficient use of bandwidth.
3.7.2 In wireless context
While wired network access mediums have well defined / bounded error rates, based essentially on the electromagnetic noise ratio, wireless mediums (such as inter vehicular communications are prone to the creation of unpredictable non-stochastic jitter and packet losses through additional issues and challenges, including [98]:
  • 1.
    Fading of the transmission: based on distance between nodes, changes in the propagation environment or node mobility.
  • 2.
    Shadowing and attenuation: from other physical obstacles and phenomena (e.g. reflection, absorption, diffraction) [105].
  • 3.
    Interference: from other devices communicating on the same spectrum frequencies. This issue is more pronounced when unlicensed spectrum (e.g. 2.4 GHz) is used since its usage is uncontrolled and prone to additional noise from other uncontrollable entities.
  • 4.
    Random channel access latency: the randomness associated with existing listen-before-talk procedures for multiple access to the wireless channels contributes to non-stochastic delays.
  • 5.
    Integration of synchronization mechanisms: while considering the differences in propagation time in wireless communications.
  • 6.
    Bounding an end-to-end latency: or its jitter by integrating scheduling on transmission by each Access Point (AP).
  • 7.
    Integration of reliability technique: while considering and taking maximum advantage of the main different channel characteristics.
  • 8.
    Resource management: including the usage of several APs for reliability or the dynamic management of the spectrum used at a given moment, which remain open challenges.
Achieving determinism within a wireless communication requires that all these challenges are properly addressed. While fading, interference, shadowing and attenuation are mostly unavoidable without changes to the physical environment, interference can be minimised. The usage of other parts of the spectrum, possibly licensed, with wider channels, and frequencies in multiple links working in conjunction (multi link operation) can further decrease the overall noise within the physical layer and therefore decrease the jitter of latencies, and increase the determinism of the communications [98]. In addition, a periodic analysis of the spectrum usage can support a decision on the best frequencies to be used.
Regarding the unpredictable time to access the wireless medium, existing listen-before-talk solutions such as Carrier-Sense Multiple Access (CSMA) consider that in the event of a collision, before re-transmission a random amount of time shall be waited. This creates an unpredictable, non-deterministic jitter on the transmission of information. Other channel access solutions must be considered in order to achieve higher determinism on the transmission of the frames through both a wired or wireless medium, such as dividing the channel per frequency, time or both (e.g. Frequency-Division Multiple Access (FDMA),  Time-Division Multiple Access (TDMA) and  Orthogonal FDMA (OFDMA), respectively) or adding scheduling algorithms to listen-before-talk solutions ( CSMA-Collision Avoidance (CSMA-CA)) [106]. In alternative, solutions that consider collisions to be solved at the receptor side (e.g.  Non-orthogonal Multiple Access (NOMA)) can also be considered.
Another promising solution is a coordination and scheduling of the access to the wireless medium to ensure not only the typical concerns of fairness or throughput but the fulfillment of a specific time of transmission, ensuring more determinism than the previous approaches [98]. Within wireless mediums, the usage of the IEEE 802.11be (Wi-Fi 7) concept of restricted Target Wake Time (rTWT) (originally introduced for power saving) can further decrease conflicts during transmission. By defining when a certain station cannot transmit — and therefore any transmission ends before the start of this time — it is possible to further decrease collisions between critical and less critical frames being transmitted simultaneously.
Projects such as PREDICT-6G [107] and DETERMINISTIC-6G [108] defend that, since the wireless environment has a non-stochastic jitter that is unavoidable (at the very least because of fading, interference, shadowing and attenuation), any system should adapt to those variations. Such adaptation can be done through online (in runtime) prediction of the variability of latency and jitter through monitoring of the system and prediction on Key Performance Indicators (e.g. instantaneous availability) to then act on resource allocation and schedules. To support this prediction, models of the channel behaviour can be considered to model the performance of the channel based on the medium conditions, transmission power, an angle or distance between nodes. Such models are not a one size fits all solution, with each different environment requiring a different analysis [98].
Regarding the integration of synchronization mechanisms, the IEEE 802.11 standards provides time synchronization capabilities through a mechanism named Timing Synchronization Function (TSF) that supports a low precision synchronization with a maximum of 1 second of accuracy [109]. The TSN standard for high precision synchronization, the IEEE 802.1AS (PTP) defines a profile for 802.11 mediums [98] that considers within the synchronization of the clocks an estimation of the propagation delay and based on the number of neighbours of the nodes.
Regarding the bounding of end-to-end latency and jitter, the integration of shaping techniques, namely TAS/802.11Qbv or ATS is possible. Solutions for integration include mapping the scheduling policies to each transmission queue within an AP and by adding to each frame fields with the flow identification and characteristics (delay bound, maximum packet size) to map the scheduling policies to the 802.11be MAC. The overall scheduling would require a full knowledge of the network topology.
Regarding reliability, including FRER (TSN standard IEEE 802.1CB) through redundant wireless links and channels is possible. The usage of several APs, with the necessary seamless predictive handover between APs, remains however an open challenge [109]. In addition, management of the transmission power can further improve the reliability by decreasing the impact of interference and obstructions in the transmission.
Finally, regarding the resource management, the need for some sort of centralization in decision making for the configuration of the network nodes — APs and switches — is supported both by TSN and DetNet standards. TSN IEEE 802.1Qcc standard define that 2 entities (the previously defined CUC and CNC) are responsible for monitoring and defining resource allocation and configurations [109], while the DetNet Reference Architecture [88] considers the existence of a Controller Plane responsible for “topology discovery and maintenance, packet route selection and instantiation or path fail-over mechanisms” [110]. An architecture should therefore consider the existence at least of a controller that configures and monitors the network nodes and the information flow.
All the previous points may benefit from additional performance performance with a deployment considering hardware acceleration, for example through the usage of Application-Specific Integrated Circuits or FPGAs [111]. FPGAs, by allowing the reconfiguration of the hardware to fit the needs of the application(s), would add to the hardware acceleration, the flexibility of deployment and reconfiguring of new wireless APs that combine the previous techniques accelerated by hardware. Such acceleration is essential to ensure the processing times are sufficiently small to allow determinism in the overall communication.
[…]
  • 1.
    Timer: used by the scheduler and synchronized by an external Pulse Per Second (PPS) signal;
  • 2.
    Scheduler: operating at OSI level 2, decides which frames should be sent and when are they due;
  • 3.
    Custom MAC: contains information for controlling frames, with changes from the original 802.11 access scheme to avoid collision between real-time and non-real-time frames and supporting the necessary synchronization between clocks;
  • 4.
    Physical layer: The encoder/decoder and additional hardware modules required for modulation features.
While this work offers a promising solution, it lacks full testing in mobility scenarios, integration with TSN and integration with other wireless standards such as 5G-NR. However, such integration with cellular communication technologies is possible. 3GPP Releases 16-18 defends that, ensuring ultra reliable and low latency communications start at the radio level and that TSN standards should be added to the overall architecture, with parts of the 5G core acting as TSN bridges [109], one per 5G User Plane Function (UPF[113]. This allows a transparent integration of 5G systems with a TSN system but require a set of translators between the 5G control plane and TSN standards [114]. The configuration of the TSN network can then be based on the 5G configuration, with the mapping between the TSN network configurations [115](such as priorities or TAS gate opening times) and 5G QoS parameters, namely the 5G QoS Identifier (5QI), a scalar that represents characteristics such as [113, 116]:
  • 1.
    Resource Type: distinguishes between Guaranteed Bit Rate (GBR), Delay Critical GBR (DC-GBR) or non-GBR, where the first two types of flows can pre-allocate resources;
  • 2.
    Priority Level: indicates a priority in scheduling resources. Can be used to prioritize the allocation of resources to certain flows;
  • 3.
    Delay Budget: an upper bound of the time a packet can spend inside a 5G system (between the User Equipment (UE) and a UPF) without being dropped;
  • 4.
    Packet Error Rate: a ratio between the number of received packets and the number of transmitted packets.
Developing algorithms to map TSN configurations to 5G configuration values remains an active area of research. The work of [117] proposed mapping the default priority value of a 5G frame to the Priority Code Point (PCP). Other work by [118] consider that TSN frames shall be considered DC-GBR, while [116] propose the mapping of TSN traffic into 5G QoS 5QI characteristics. For example, the resource type characteristic is derived from the type of TSN flow: deadline constrained, jitter constrained or bandwidth constrained, per the truth table in table 7, with similar analysis for the other characteristics within 5QI.
Table 7: Mapping between TSN flow characteristics (on the left) and 5QI values (on the right)  [116].
Deadline Jitter Bandwith GBR DC-GBR Non-GBR
0 0 X 0 0 1
0 1 0 1 0 0
0 1 1 0 1 0
1 0 0 1 0 0
1 0 1 0 1 0
1 1 0 1 0 0
1 1 1 0 1 0
[…]
3.7.3 In vehicular context
Solutions to adopt TSNs within the vehicle’s systems are sparse and moreless recent, with some of the oldest integration within vehicular systems in 2016 [120], and further integration in recent years [121, 122], including with 5G [123, 124]. In addition to 5G, both wired and wireless technologies based on IEEE 802.11p or IEEE 802.11bd should be considered within these networks, ensuring that service(s) can communicate with the vehicle’s systems and surroundings with reliability and redundancy.
The challenges on TSNs within vehicular contexts are two-fold: for one, the automotive use case is a critical scenario, which requires well defined latencies, ordering and high level reliability; on the other side, the high dynamism of the intra and inter vehicular topology — each vehicle may have a different set of sensors and actuators, creates the need for a more automatized TSN configuration that can be adapted to different streams.
The work of [125] discusses how to characterize the traffic — frame size, period, deadline constrains, priority, proportion of streams and how to analyze efficiency — overload analysis, scheduling analysis, and spacial complexity/memory usage. This can be a basis for a TSN configurator to decide the mapping between the streams within the vehicular network and the necessary TSN shapers configurations.
Not only TSN has to be properly configured, but interoperability with other protocols must be assured. Ethernet with TSN must coexist and be interoperable between with other wired and wireless access technologies, including legacy automotive standards such as CAN. Synchronization strategies between the vehicular Ethernet backbone and CAN are currently being investigated by [14, 67], with three approaches under consideration as depicted in fig. 13:
  • 1.
    Periodic snapshot: periodically all original frames are mapped unto 1 or more frames of the destination bus. Within this solution, the main challenge is the computation of the period of snapshot, specially in asynchronous traffic.
  • 2.
[…]
  • 3.
    All packing: in this solution, all CAN frames up to a point (functionally defined) are grouped into 1 Ethernet frame. This solution optimizes the usage of MTU of Ethernet vs the 1:1 packing solution.
Figure 13: Examples of CAN from and to Ethernet Gatewaying Strategies [14].
The integration of deterministic communications in vehicles is not exclusive to TSN standards. DetNet standards can be used in conjunction with TSN standards to provide both upper and lower latency boundaries (while TSN can only ensure an upper bound), better scalability and interoperability with existing IP infrastructure (since it acts on OSI layer 3) [126]. Both sets of standards can act simultaneously, with [126] proposing that each TSN domain would be mapped into a subnet (or a tunnel) of a DetNet network, as depicted in fig. 14.
Figure 14: Examples of integration of TSN and DetNet may be considered within a 5G system [126].
Integration with cloud infrastructure is also possible. For example, within [127] is proposed a full-fledged dynamic interference-aware cloud scheduling architecture for latency-sensitive workloads.
4 Open challenges
The previous subsections focused on specific parts of the vehicular stack - either middleware, Operating System and virtualization techniques. All these parts need to be configured properly in a holistic way to ensure the services are deployed properly. The work of [30] emphasizes the need for configurations and reconfigurations of the whole vehicular stack to be dynamic, from the subjacent bus (such as Ethernet), up to the configuration of the ECUs and services. Such reconfigurations require the integration of multiple levels of configuration — from deterministic networking standards to the appropriated SOA protocols to higher level as described in fig. 15.
Figure 15: OSI layers to have into consideration for an automated configuration of the whole vehicular stack [30].
Per the current state-of-the-art, it is essential to create tools to support a more agile service development, that considers the services requirements in terms of bandwidth, safety, criticality and timing of delivery of its information, while ensuring a interoperability and modularity to support (almost) any vehicle. These tools shall take into account not only the orchestration of those services throughout the vehicular zonal E/E architecture and its ECUs but also the network interconnecting those ECUs and connecting the vehicle to other elements on the smart city road. Therefore, determinism communications and its standards, tendentially and historically thought for static well-known environments, need to be able to be dynamically configured to support the highly dynamic environment within vehicles and their equally dynamic surroundings. Other challenges include supporting and improving:
Figure 16: Overview of the potential use cases.
  1. 1.
    Agility of service development: simplify and harmonize vehicular development processes — (almost) everything as software and software processes;
  2. 2.
    Data driven services: vehicular architectures should be data driven in order to process great quantities of data from camera/lidar/sensors streams in use cases like autonomous driving [128];
  3. 3.
    Safety concerns: following zero trust policies is essential, as well confidential computing concerns.
  4. 4.
    Optimized resource orchestration: leveraging the unused resources — but not necessarily on a in-vehicle point of view, while considering costs and capabilities (CPU bound, I/O bound) from the hardware into the scheduling and orchestration;
  5. 5.
    Interoperability: multi-vendor, multi-hardware and multi vehicle interoperability;
  6. 6.
    Modularity: towards a more manageable stack where automakers don’t do everything;
  7. 7.
    Next-gen distributed storage: considering cloud storage and solutions such as InterPlanetary File System (IPFS) towards use cases such as OTA updates;
  8. 8.
    Clustering: support the clustering of ECUs and the deployment of vehicular services across clusters, with reinforcement learning techniques (or equivalent) to choose the best node affinity;
  9. 9.
    Automatic generation of design architectures: computer generated architectures for zonal vehicular ECUs;
  10. 10.
    Legacy compatibility: solutions to integrate ECUs interconnected with deterministic communications in a zonal SOA architecture;
  11. 11.
    Virtualization and automatic scheduling: reallocate workloads depending of resources with general purpose hardware and virtualized infrastructure;
  12. 12.
    Digital twin: standardize the vehicle state representation [30].
Considering the current state-of-the-art, within the dynamic deployment of vehicular services focusing on the communication part with TSN and deterministic networks, the main challenges include:
  1. 1.
    Easy to use/plug and play: Auto-configuration of parameters for TSN shapers based on the number, characteristics of ECUs and workloads. For example  [55] propose a TSN platform abstraction layer and set of API for applications to send/receive frames. The API was responsible for managing flow registration, transmit and receive Ethernet frames, creating the Ethernet frames, configuring the Ethernet port and scheduling frames to the appropriate queue.
  2. 2.
    Integration of TSN within wireless networks: Some efforts are being done, but it is current work as discussed in section 3.7.2 [129].
  3. 3.
    Built-in time awareness: at the link layer — at the applicational level is more resource intensive, while at link layer could even be accelerated by FPGAs.
  4. 4.
    TSN shaper automatic selection and configuration: automatic generation of necessary configurations, with reinforcement learning as a possibility to choose the best configurations.
The next sections will focus on discussing potential use cases that may benefit from such solution and strategies, while also discussing the proposed solution for a more agile service development, and tools for its evaluation.
5 Potential use cases
Figure 16 lays out the main areas of use cases where the SDV paradigm may be useful.
Even if vehicles will tend to become driver-less, autonomous vehicles may also take decades to be fully on the roads, and they will coexist with manual-driven cars. As such, autonomous vehicles shall consider the needs of the other occupants. It is therefore essential to think about vehicles, not only as means of transportation that may tend towards not requiring drivers, but also as adaptive objects, that suit themselves to the needs of occupants.
Examples of such needs would be the vehicle itself recognizing the needs or potential health issues, and providing personalized profiles and automatic adjustments on seating, cooling systems, driving profiles, infotainment screens, lighting, audio, and other vehicular subsystems that may support some sort of impairment. But this adaptation must be carried out flawlessly, since any failure can mean harm to someone, or even death.
Table 8: Main requirements for some possible use cases (based on [130], with adaptations).
[…]
  • a
    ETSI CAM size with extras
  • b
    considering maximum velocity of 120 km/h
  • c
    maximum non highway speed of 90 km/h
  • d
    urgent, during a vehicle top speed at certain highways
  • e
    during a vehicle top speed at certain highways
  • f
    heavy traffic around the vehicle
  • g
    worst case scenario, with human reaction
  • h
    majorant of the provided values
  • i
    speed, location and additional information
  • j
    minorant of the provided values
  • k
    highway / rural / urban
  • l
    maximum velocity of 120 km/h
  • m
    considering an urban situation
  • n
    worst case scenario in possible range of frequency of transmission of Manoeuvre Coordination Message (MCM), per [131]
  • o
    maximum size of the message [131]
When considering areas of use cases for SDV, one can consider both the drivers — for which most use cases, including autonomous driving, tend to focus on — and occupants. Both can benefit from the vehicular technology that adapts to each occupant while increasing its autonomy. Software-Defined Vehicle can use the existing sensors and actuators on vehicles to support new services to support drivers and occupants in situations automakers could not predict during their design phase.
The following subsections present some use cases that are envisioned as possibilities to validate the system(s) to be researched — intra and inter-vehicle deterministic use cases and critical non-active safety use cases that require adaptability — both at software level and network level (while ensuring reliability and bounded latency). In all described use cases, the SDV paradigm supports some adaptability of the vehicle to the occupants’ characteristics and state to support some automated action (instead of a manual process). The occupants, either driving or passengers, can greatly benefit from services for customization of vehicular functions, protection, monitoring, and companionship.
5.1 Requirements and characterization of vehicular services
Table 8 provides an overview of the main requirements envisioned for the use cases. Use cases can be characterized by several types of requirements. The study in [132] defines several vehicular use cases and their characteristics in terms of deadline (soft/hard and interval), data complexity (simple/complex and large) and pattern (discrete/continuous).
During the overall definition and characterization of any vehicular service supporting an use case, it is essential to formulate the service and characterize its tasks in terms of those requirements, with particular emphasis on characteristics such as [19, 39, 133, 134]:
  • 1.
    Priority and criticality;
  • 2.
    Data specifications: such as format, size and real time guarantee requirements (periodicity or deadline for receiving/sending data);
  • 3.
    Expected computing location: edge vs cloud;
  • 4.
    Computing resources: either in terms of CPU, memory, storage, or graphical resources;
  • 5.
    Required real-time guarantee levels: in hard real-time, weakly hard real-time or soft real-time, depending on the impact of missed deadlines;
  • 6.
    Characteristics specific to the vehicular domain: such as the mobility of vehicular nodes and associated communication handovers and failures, and the heterogeneity and overall low specification of the ECUs;
  • 7.
    Non-functional requirements: noteworthy examples include high interoperability (including backwards compatibility with older vehicles and other vehicle brands), reliability, availability, scalability, extensibility, adaptability, testability and cost minimization.
These characteristics can be seen as the Service Level Agreement (SLA) of the service. Such SLAs should and could be defined by the developer of each vehicular service within a descriptor file for the service to support selecting the best technologies/stacks and configurations. Such configurations would be ensured/enforced by a distributed configurator through a set of QoS tools to enforce policies on the task scheduling and communication technologies tasks [134]. On task scheduling, the configurator would be responsible for a dynamic, intelligent, task scheduling and resource estimation, provisioning and allocation, including the offloading to other layers of communication (fog and cloud) while keeping the original non-functional requirements. Such scheduling could include the prediction of the overall resource usage while considering heterogeneity and computational cost of the nodes for optimization purposes [135, 136]. Estimation and prediction of the mobility of a certain service deployment based on trends of resource usage and prediction of the mobility of the vehicle would be essential in scenarios where the service requires deployment on multiple vehicles and/or external infrastructure to ensure coordination with other vehicles or entities [137].
Within the communication technologies, the configurator would be responsible for the dynamic management of the communication technologies during the mobility of the vehicle. Solutions would include the federation of multiple physical wireless and wired access layers in order to obtain a multi technology solution. Such management could also include slicing of the communication solution to support different levels of QoS specified by the service descriptors, as well as the management of the high mobility with the management of reactive and proactive handovers leveraging SDN [138] based solutions;
5.2 Services for autonomous driving in all road situations
While autonomous driving is a current and future trend with great focus from automakers, its adaptability to all road and traffic situations is an open problem. Current solutions can handle open road or highway contexts, but higher density (e.g. urban) contexts require a more cooperative approach to autonomous driving, with vehicles sharing all the information they possess and coordinating manoeuvres.
5.2.1 Context Aware Maneuvering
Considering the importance of working towards autonomous driving for everyone, it is vital to ensure that each autonomous driving vehicle is capable of creating the best possible decision not only for itself, but for other vehicles, and road occupants as well. This requires leveraging the concept of cooperation between vehicles and others. While standardization work, for example through European Telecommunications Standards Institute (ETSI) MCMs are under way, the cooperation between vehicles, particularly for autonomous driving is in its initial stages.
Vehicles may consider cooperative manoeuvres between non-autonomous vehicles and/or autonomous vehicles to ensure that complex manoeuvres are done with no harm. For example, within the context of the cooperation between a personal vehicle public and private transportation, a service can be considered to give priority to bus, ambulance or other public transportation to enter a certain crossroad instead of a privately owned vehicle. This service shall be react to changes on the environment and surroundings of the vehicle and decide the action of each vehicle. Both receive information on the road situation and sending the proposed action to other vehicles would require a minimized bounded latency to ensure a proper decision and that such decision is enforced in the right moment.
When considering the air mobility paradigm, in particular electrical Vertical Take Off and Landing (eVTOL) vehicles, the cooperation between the vehicle and other vehicles gain an additional importance. New scenarios shall be considered, including cooperation between vehicles and the eVTOL and the road infrastructure (such as landing pads) to ensure that no accident occurs during take off or landing, even in bad visibility (by occlusion or lack of lighting). In these scenarios, coordination between vehicles and infrastructure can avoid a catastrophic air accident between the vehicles or a potentially deadly crash on top of property or people [7].
Autoware, an open source framework considered to be state-of-the-art for autonomous driving [^53] Apart from closed source solutions from automakers (more details on section 2.1), where it is not possible to fully understand the scope of the cooperation between vehicles and its interoperability with other brands., considers several services to ensure an autonomous driving — albeit in previously mapped environments — to support manoeuvres such as curb-to-curb, lane following, lane merging, parking or reaction to traffic lights and dynamic entities (e.g. Vulnerable Road Users).
To produce plans that eventually become decisions and actuation to the engines, braking and other vehicular subsystems, Autoware relies on information of the vehicle itself and on a model of the environment. All this planning is therefore very local, with cooperation from other vehicles and road entities proving essential not only to improve Autoware’s vision of the world —in a cooperative sensing solution such as the one defined by ETSI Collective Perception Service (CPS)—, but also to coordinate the manoeuvre that each vehicle will do, ensuring that a safe manoeuvre for one vehicle does not possess danger to another.
When the planning becomes cooperative and not only local, a new problem arises. Information from and to each vehicle must be sent within a well defined time-frame that is compatible with Autoware producing a decision and actuating on the vehicle’s subsystems. If the information arrives after this defined time-frame, the decision will not be correct, while if it arrives before this defined time-frame, the decision might be unnecessary — but also potentially dangerous [^54] For example, a vehicle would end up braking before being necessary, becoming vulnerable to being hit in the back..
As an example scenario, one may consider coordinating lane merging manoeuvre between an ambulance and a private vehicle. This use case requires the development of services depicted in fig. 17, namely a) a message creator and producer to support exchanging data between Autoware instances in different vehicles, and b) a parser and injector of such information into Autoware planning. The data format would follow ETSI MCM standards and it would be periodic, with deadline insurance requirements (and therefore both a minimum and bounded latency requirement) with no tolerance to failures of delivery, and with a variable size with an upper bounds corresponding to the maximum size of the MCM message.
Figure 17: Overview of the Context Aware Maneuvering use case.
The integration of Autoware with cooperative manoeuvering is considered in [139] [^55] Additional details at tlab-wide.github.io/AutowareV2X.. However, it fails to consider the deterministic communications requirements or the need for orchestrating the overall deployment in vehicles’ ECUs in a more automated approach.
5.2.2 Vehicle malfunction situations
Even when full autonomy of a vehicle is not only a reality but also legal, vehicles can still have unexpected issues that endanger their occupants and surroundings. Examples include a flat tire, a tire explosion, failed waterproofing of the batteries (in the case of electric vehicles), or a failure on the valve timing (in the case of internal combustion engines). In these scenarios, services to analyze, monitor and warn others about the situation should be deployed in a bounded minimal time, and the information to other vehicles shall be sent with both minimized, and bounded latency.
When considering Connected Autonomous Vehicles, services that monitor and react when sensors are malfunctioning are critical, and their information must also be bounded in time and transmitted reliably. Without such services, the CAVs cannot react to a malfunctioning or occluded sensor (such as camera, LIght Detection And Ranging (Lidar) or equivalent) and can wrongly assume that their information is correct, with potentially deadly consequences.
5.3 Services for vehicles tailored for everyone
While the future of mobility is seemingly more autonomous and more feature-packed, less concern is given to occupants and on their needs, including children or occupants who may have some sort of physical or cognitive restriction. Restrictions considered here may include some minor impairment, vision or hearing issues, concentration issues, and others, that restrict the occupants’ capability of reacting to the environment in the same way.
The following situations are examples where a vehicle has an occupant with some issue/limitation who can greatly benefit from a SDV paradigm. In this paradigm, the vehicle can deploy services tailored for the people’s specific needs for customization of vehicular functions, protection, monitoring, and companionship. In all situations, like in previous use cases, the services to be developed and deployed shall have their data delivered in a bounded time — so the services are not left in an inconsistent state.
5.3.1 Automatic vehicle configuration based on occupants characteristics
Information on the occupants’ characteristics and vital parameters can be obtained by intra-vehicular subsystems that allow the vehicle to adapt interior functions (e.g. seating position and configurations, lighting, sound, infotainment) to the best comfort of the occupants, therefore creating a set of customizations tailored to each occupant’s characteristics.
Not only the services can adapt, but warn if there is an issue (like a non-fixed wheelchair or baby seat). Such warnings can be sent both to the infotainment vehicular subsystems, and also to other vehicles. Vehicles can then use such information, as well as information about the occupants’ characteristics, to deploy services to adapt the vehicle driving mode, profile and characteristics. For example, such adaptation could signify an increase cruise control following distances, reduce throttle response, or warn other smart city infrastructures when someone inside the vehicle is ill or impaired.
These services can greatly benefit the healthcare industry in particular. Consider a scenario where an ambulance is transporting a patient. Vehicular services can support scenarios such as this:
  • 1.
    Ambulance: I have a person in risk of life.
  • 2.
    Other vehicles: Understood, I need to slow down and provide space around the ambulance.
  • 3.
    SDV Orchestrator: We’ll provide you with specific applications (e.g. to control acceleration profiles or climate control settings). We’ll deploy an application for monitoring ill people that will run on a cluster of $X$ vehicles. We will coordinate with all the vehicles in the path, and with the city traffic manager, to create a corridor for you to be able to pass between other vehicles in the city.
5.3.2 Adapting the vehicle’s subsystems to impaired occupants
Consider a situation where, in order to transport impaired people in a wheelchair, a healthcare provider must [^56] From an informal interview with a specialist nurse in medical-surgical nursing.:
  1. 1.
    Get the wheel chair up a moving ramp or lift, per depicted in footnote 58575757Another example of such ramps and lifts in gilaniengineering.com.au..
  2. 2.
    Attach the wheelchair to the vehicle through a set of several belts fixed to specific points within the vehicle, as depicted in footnote 59.
(a) Example of ramp where wheelchairs need to be pushed manually by the healthcare provider (courtesy of Tripod Mobility [^58] More details at tripodmobility.com.).
(b) Example of expected attachment points (4 in this case) within a vehicle for a wheelchair (courtesy of BraunAbility [^59] Available at braunability.eu.).
Figure 18: Detail view on the main steps for transporting impaired people on a vehicle.
The healthcare provider must do such manually, multiple times (for multiple people) and in a short time-frame and without little to no feedback on whether the person and the chair are actually fixed properly. Mistakes include fixing outside of the fixing point(s), twisting the straps and/or simply not fixing the person and/or chair. In addition to that, specific formation is required to do such tasks.
A vehicular service (or set of services) that is capable of communicating with other in-vehicle systems (e.g. infotainment), but also with the healthcare provider registry systems, would greatly decrease the potential human errors of the current solution, and since it would minimize the dependency on humans, it could even support an increase on the patient’s privacy. One could even consider a more intelligent connected wheelchair directly coordinating with the vehicle all the necessary movements. Such systems would not require a minimized latency, but guarantees of delivery of the information within well-known time bounds and with low jitter in order to ensure no unwanted movements of the wheelchair that may harm the person.
The adaptability of the vehicle does not need to end on improving or automate mechanical subsystems. In fact, greater challenges can be considered. Depending on the profile of the person (whether it’s any sort of impairment, special needs or old age), the vehicle could further adapt itself for example by choosing a softer, appropriated music (supporting music theraphy [^60] More details at britannica.com/topic/music-therapy.), adapt the climate subsystems to ensure the adequate temperature, tint the glass to avoid too much sun exposure.
Other scenarios may include the protection of other type of vulnerable humans, specifically babies:
  • 1.
    Managing situations were babies are left unattended in cars. It is a widely known issue [^61] Unfortunate examples at cnn.com.. Babies, when left in vehicles and especially in hot weather, where vehicle temperatures can rise rapidly, are at risk of heatstroke and potential death. Vehicular subsystems and services could detect the presence of the baby and alert caregivers or authorities if the child is left inside under dangerous conditions.
  • 2.
    Managing the set of airbags that should or should not be deployed based on the position or age of the baby.
  • 3.
    Monitoring a baby who is ill or very restless and provide such information directly to a hospital.
5.3.3 Personal assistants
Vehicles can include humans in new ways within the driving experience, even when considering the SDV paradigm is parallel to the advent of autonomous vehicles. The unparallel amount of information and sensing the vehicle can obtain from its environment, driver and occupants, can be used to support personal assistants that give occupants the sensation of being in control — even if the vehicle is self-driving, potentially increasing not only the enjoyment of the vehicle but more importantly the perceived trustworthiness on the vehicle.
Such assistants can be developed as a set of subscribing services within the vehicle’s ECUs dedicated to infotainment subsystems and, if using speech or touch interaction modalities, it can greatly benefit from a well defined timely delivery of information which decreases the overall jitter visible to the end users.
5.4 Services for assisted fleet management and maintenance
SDVs can also support scenarios related to management of vehicle(s), including managing its health, maintenance (predicatively or not) and homologation, both for individual vehicles or for a fleet.
5.4.1 Fleet management
Fleet management scenarios and remote services, such as the one presented at fig. 19626262More details at autowiz.in., are examples of use cases that can benefit from a set of services deployed across multiple vehicles. They consider a set of services containing the logic required to monitor, in real-time, the status (e.g. location or number of hours the vehicle is driven) of a vehicle within a fleet. This set of services, if orchestrated throughout a cluster of vehicles, can support systems for enterprises managing a vehicular fleet, homologating each vehicle, or even for state entities to perform the routine inspection. The information from/to those services, while it does not necessarily require an extremely low latency, it must be delivered and delivered within known time bounds and with low jitter.
Figure 19: Example of vehicle fleet management system (courtesy of AutoWiz).
5.4.2 Remote vehicle health and predictive maintenance
Constant monitoring of the main vital parameters of a vehicle state — such as tire condition, brake system condition, engine cooling system status (in the case of internal combustion engines) or the power delivery subsystems (in the case of electric vehicles) — requires not only a reduced latency but also, in case the vehicle detects an issue, a guarantee of delivery in a bounded time. This is required so that, at the very least, other vehicles are aware of the problem before it has impact on the road. In the case of emergency scenarios — as previously described in section 5.2.2, the latency shall also be minimized.
Predictive maintenance scenarios, while not necessarily requiring a bounded latency within its communication, still require a dynamic deployment of services that considers the vehicular ECUs.
5.4.3 Over-the-air updates and real-time homologation
Changes in the vehicles’ software — either via OTA updates, through the addition of new services or after a vehicle is sold to ensure no changes — should, in an automated way, be reflected in a re-homologation process. This process, while not requiring a minimized latency, shall be done in a bounded timely manner, with a guarantee of reliability; otherwise, vehicles may be driving on public roads with an unwanted, unsafe state [140].
6 Proposed vision
The SDV concept is capable of supporting several services spanning several different types of use cases. All the previously defined use cases require $N$ time critical services and $K$ services. In all these use cases, a certain vehicular service can either be fully deployed on one or more nodes, or divided through multiple nodes / cluster of vehicles, with both scenarios considering the cloud as an option. Also common to all use cases is their need for time criticality. Within the SDV paradigm, the vehicle may send the information about its occupant to other subsystems, or vehicles through communications, which must be deterministic in order to ensure the information is obtained not too soon nor too late, therefore risking lives. While the use cases do not necessarily focus on controlling the critical systems intra-vehicle such as power-train and braking subsystems, they are essential for other passive safety functionalities.
Given the criticality on the one side of the use cases and associated services, and the increased complexity of development and deployment of services, it is essential to consider a dynamic management of services. Towards such management, several layers must be considered as depicted in fig. 20. These layers should act at least in the following domains [30]:
  • 1.
    Orchestration of the services: where they should be run and policies to decide the node(s) and the number of replicas. Such domain is the responsibility of the orchestrator, a .
  • 2.
    Configuration of the abstraction layers required to support exchanging data in predefined formats and protocols. Such domain is the responsibility of the data layer configuration ( b ).
  • 3.
    Management of the deterministic network: including automatic configuration of routing and deterministic traffic shapers and reliability processes. This domain is the responsibility of the deterministic network configurator ( c ).
  • 4.
    Virtualization of the hardware in order to support better resource usage and decoupling of services and operating systems from each ECU. This domain is the responsibility of the vehicle abstraction layer ( d ).
Figure 20: Overall layered view of the architecture main modules framed within the OSI model.
Figure 21 presents the overview of an initial proposal of architecture towards such a dynamic management of vehicular services. This architecture supports support service scheduling, redundancy, and healing based on the vehicles’ computing capabilities.
Figure 21: Overall high level architecture of the proposed solution, depicting the service development and deployment flow from the service developers down to the vehicular hardware.
The Controller ( a ) is the main entry-point of the system to be researched and ensures the first aforementioned layer. Through user-defined and system-defined metrics within a Service Descriptor, it decides the amount of resources and the placement of those resources throughout the computing continuum and the different vehicles throughout the smart city, while considering possible handovers and optimal usage of the computing and network resources. The controller is responsible for controlling 4 verticals.
The first vertical, the Deterministic Network ( b ), ensures that messages – either control messages to actuate or status messages containing information from vehicles and the smart city – are delivered in a well-specified time bound and with fault-proof reliability. Research topics here include, but are not limited to, integrating clock synchronization mechanisms into wireless vehicular networks and mechanisms towards dynamic automatic definition, and optimal configuration of flows of packets, their queue priorities and characteristics. The module within the architecture, the Deterministic Network Configurator, is responsible for creating the necessary traffic flows, their parameters and configuring the necessary hardware, namely the Network Interface Controller/Card (NIC) within each ECU and the TSN capable Ethernet switches interconnecting them. It is expected at least 3 traffic classes to be configured: a) system control messages with maximum criticality; b) high size (non) periodic streams of data and c) periodic or non periodic service messages.
The Data Layer vertical ( c ) is responsible for ensuring the data is exchanged between the vehicular services in the most appropriated set of SOA protocols — such as SOME/IP, MQTT or DDS. Research topics here include, but are not limited to, the design of a abstraction layer. Such abstraction layer would increase interoperability between all these protocols while ensuring deterministic communication through integration with the underlying deterministic network mechanisms.
The Orchestrator vertical ( d ) is responsible for ensuring the deployment of the services, either within a Kubernetes worker node, within a virtual machine managed for example by KVM or through a containerd or equivalent container runtime. Research topics here include ensuring deterministic communication through integration with the underlying deterministic network mechanisms (as previously mentioned in the Data Layer vertical).
These 3 verticals dispatch orders to the Vehicle Abstraction Layer (VAL) ( e ) of each ECU, which abstracts different types and models of vehicles’ ECUs – both simulators and real vehicles – their actuators, sensors, and ECUs through the abstraction of existing frameworks, such as AUTOSAR, Autoware or others. Conceptually, the VAL is no different to the standardized abstractions operating systems created during the mid 1980s and 1990s to abstract the hardware from the remaining kernel, with [133] proposing an extension of such a paradigm into the vehicular paradigm, as depicted in fig. 22.
Figure 22: The VAL based architecture for vehicles,as described by [133], where the divisions between trusted and untrusted zones and between safety and non-safety services are evident
The research on the VAL may also consider the creation of a digital twin of the vehicle and the research of the impact of sensors and services on the vehicle networks.
6.1 Overall end-to-end interaction
Figure 23 provides an overview of the end-to-end interaction of the proposed system, covering everything from the features available to the end-user to the configuration of the required hardware needed to run the services and meet their requirements.
Figure 23: Overall end-to-end interaction model of the proposed solution.
Our vision includes a system that manages service deployment, execution and evaluation of intra- and inter- vehicular services. Services can be associated with domains, that define not only requirements but the universe of information the service can subscribe. Service developers define a service descriptor (an example of the current vision for its structure is detailed in listing LABEL:listing:arch-use-case-service_descriptor) to describe the main service characteristics. This set of requirements is used to configure and dispatch the services to the appropriated orchestrator in the appropriated ECUs and configure the datapath, i.e. the network and data layer resources required for the service to communicate with others.
Services must be deployed by the controller which must decide the best node(s) for the service characteristics, either automatically, manually, or based on criteria (e.g. time, energy, cost). The controller must further ensure latency bounds for the deployment of the service and for the exchange of messages between them if the service descriptor requires criticality and/or QoS levels on the data exchange and processing. The controller is the module responsible for a) evaluating the nodes on the match between the service characteristics and the node characteristics, and b) dispatching the configurations to the $N$ nodes best fulfilling the service characteristics. Receiving as input both the service descriptor and the node characteristics, the algorithm outputs a score to each node and if the node may be used or not (to support requirements that must be fulfilled).
The vehicular service developer shall interact with the system through a declarative paradigm, i.e. set the wanted state [^63] Kubernetes is a widely known example of such declarative paradigm.. A service descriptor file — or equivalent structure — describing the necessary characteristics of the vehicular service can support this paradigm. This information is defined by the service developer and can be used by the system to be developed to support the service deployment.
Associated with each field is not only the value, with the type prescribed, but also its criticality. Since not all conditions can be met by a node, it is essential for the controller module to know which metrics can be unmet and by what margin.
For example, when considering the use cases in section 5.3.2, one can imagine the service developer creating a service descriptor such as the one in listing LABEL:listing:arch-use-case-service_descriptor. This service descriptor is then submitted to the controller within the target vehicle(s). Within each vehicle a controller is responsible fo evaluating the ECUs and configuring them and the necessary network and computing components to ensure that the service requirements are properly considered. In the case of the considered service, this entails at least:
  • 1.
    Deploying the necessary containers;
  • 2.
    Allocating and configuring the necessary resources within an ECU connected to the necessary actuators (in this case the motors for the fixing points and the ramp);
  • 3.
    Configuring the switches and other network equipment connecting such actuators and the ECUs where the service connect with the necessary deterministic networking mechanisms (in this case TSN TAS scheduling could be a possibility).
while keeping fulfilling the other already running services and their requirements.
Listing 1: Example of possible service descriptor.
title = "WheelchairDriver"
[ServiceMetadata]
Author = "TheWheelChairCompany"
Version = "1.0"
Domain = "safety"
[Flows]
[Flows.Flow1]
[Flows.Flow1.NodeSpecs]
[Flows.Flow1.NodeSpecs.NodeA]
Image = "thewheelchairservice"
ImageType = "docker"
Replicas = 2
CPU = 2
Memory = 1024
Storage = 51200000000
GPU = false
Energy = 1
Offloading = false
[Flows.Flow1.DataSpecs]
DataFormat = "json"
DataSize = 8096
[Flows.Flow1.TrafficSpecs]
Guarantee = 4
Reliability = true
Delivery = true
Wired = true
[Flows.Flow1.TrafficSpecs.TrafficTimeSpecs]
MaxLatency = 50
Periodicity = 100
TransmitOffset = 0
Jitter = 10
6.2 Proposal of an evaluation architecture and metrics
The evaluation of the proposed system requires at least a testbed with a topology equivalent to the one depicted in fig. 24. In this topology, a vehicle form a deterministic network domain and can connect to other vehicles and smart city infrastructure. A smart city infrastructure, including its cloud and RSUs form another deterministic network domain [85]. Each domain is responsible for ensuring a deterministic end-to-end latency for messages of vehicular services by configuring appropriately the necessary traffic shapers and synchronization between TSN capable switches. These switches are connected through Ethernet and ensure connectivity between the ECUs (in the vehicle domain) or between RSUs, and between RSUs and the smart city cloud (in the smart city cloud).
Figure 24: Proposed testbed solution for evaluation of the system to be developed.
In terms of required equipment, such a testbed shall include at least one or more ECUs, a data center [141], a vehicular edge computing area [141], one vehicular sensor, one vehicular actuator, a Ethernet backbone with TSN capable switch, and a CAN legacy bus. The testbed may be deployed either by software or hardware. Software simulators can be used to mimic: a) the network, b) the vehicular topology, and c) the inter vehicular domain, while hardware can be used to function/emulate ECUs, Ethernet switches with TSN support, CAN based sensors and actuators.
Analysis of overall performance of the system should be at least in terms of throughput and latency for edge (intra-vehicular), cloud, and hybrid approaches but also consider additional metrics, such as, inter-frame/packet latency [142], data rate, transmission queue length [143], percentage of missed deadline or offset of period and worst frame/packet latency. In terms of metrics it can include average node (ECU) pressure, uptime and downtime of each service (and associated error) and time until recovery.
6.2.1 Hardware equipment
The hardware equipment that may be considered within the testbed in fig. 24 shall consider hardware equipment in two different contexts: the network infrastructure, comprised by the switches, routers and access points that shall be compatible with deterministic communication standards, such as TSN standards; and the Intelligent Transport System (ITS) intra vehicle hardware, including generic computers or specific computers (i.e. ECUs).
A basic topology to simulate an entire vehicle is depicted in fig. 25 and requires at least two computational nodes, one per vehicle, connected through a network backbone (initially considered as a Ethernet Switch with support for deterministic communications), one sensor, one actuator, and necessary transceivers to bridge the sensors and the actuators to the Ethernet/Wireless domain.
Figure 25: Basic topology to simulate the vehicle.
Regarding the network infrastructure, solutions from closed source from typical network manufacturers to open source solutions are available. In principle all equipment supporting the basic architecture within [144, 145] are enough to support the basic testing of the proposed architecture. Cisco possesses TSN capable switches, such as the IE 4000 Series Industrial Switch. Within its operating system, the IOS, it is possible to install support to TSN standards [^64] Details at cisco.com..
Closed source alternatives include for example the NXP SJA1105TEL [^65] Details at nxp.com. switch, specifically target at vehicular ECUs, or the Broadcom BCM53570 System-on-chip (SoC).
Relyum offers closed-sourced FPGA based alternative solutions [^66] More details at relyum.com. for TSN switches, with theoretical support for configuration through NETCONF, with alternatives from SoC-e [^67] Such as the alternative at soc-e.com. also available.
In addition, several SoC alternatives also support TSN. Xilinx SoC, such as the KR260 series [^68] Details at xilinx.com. contain support for TSN, integrate a FPGA that may support the hardware acceleration of parts of the proposed architecture in fig. 21 and support the interconnection with existing legacy vehicular systems through its support for CAN bus, while also keeping an open architecture having in mind the reconfigurability that FPGAs support.
Regarding vehicular hardware, two alternatives may be considered: an approach based on simulating parts of the vehicle; and an entire development vehicle.
PixKit [^69] Details at pixmoving-moveit.github.io. is a possibility currently under study. Alternatives such as the The Car Lab’s HYDRAONE [^70] Available at weisongshi.org/hydraone. could also be considered, even if less advanced than the PixKit alternative, in particular without out-of-the-box semi autonomous driving [^71] Semi autonomous since the PixKit solution is fully dependent on previously mapped environments.. Within the vehicle, computation systems such as NXP, through its BlueBox Development Kit [^72] Details at nxp.com. or Renesas through its R-Car H3e-2G Starter Kit [^73] Details at renesas.com. can be considered as ECU for testing purposes. Well known alternatives such as the Raspberry Pi and equivalents can also be considered.
It is important to point out the usage of existing infrastructure, such as Aveiro Tech City Living Lab (ATCLL) smart city infrastructure — namely RSUs — to ensure connectivity of the testbed vehicular systems with the remaining smart city and other vehicles.
Finally, one should consider hardware equipment related to the use cases themselves, namely sensors and actuators. For example, the use case in section 5.3.2) includes step motors — for the ramp or fixing the seat belts — or sensors to detect unbuckled seat belts, weight sensors to detect used wheelchairs and screens for the direct interaction with the users.
6.2.2 Software and simulators
When hardware equipment is not available or if it does not correspond to the requirements of the testbed as depicted, software simulation alternatives may be considered.
While solutions such as AUTOSAR simulate and model the automotive Ethernet, they lack TSN support for such modelation [2]. To simulate the network, both in its wired and wireless components, one can use use case agnostic solutions such as OMNET [^74] At omnetpp.org., that with INET [^75] Available at inet.omnetpp.org. and with CoRE4INET [^76] Albeit with notable limitations, such as no simulation of frame preemption or clock synchronization protocols, with TSimNet model being an alternative for both frame preemption and replication and elimination [2]. or NeSTiNg [^77] Most logic related to NeSTiNg is now integrated within INET. can be used to simulate networks with TTEthernet and TSN protocols [146].
OMNET can also be extended with SUMO [^78] Available at eclipse.dev/sumo. through frameworks such as Veins [^79] At veins.car2x.org., an open source framework for running vehicular network simulations [147]. It can also be extended for service-oriented communication, through the SOA4CoRE [^80] Service-Oriented Architecture for Communication over Realtime Ethernet, at github.com/CoRE-RG/SOA4CoRE. initiative. Mininet, a popular alternative to create virtual networks with complex typologies and greatly used in evaluation for SDN contexts, can also be extended to consider wireless communications through the Mininet-WiFi [^81] At mininet-wifi.github.io. initiative and can therefore also be an alternative situation.
Matlab, within its Vehicle Network Toolbox, also allows the modeling of vehicle networks. Middleware layers, such as DDS can be integrated within MATLAB and Simulink tools as a block-set [148]
RealTime-At-Work Pegase [^82] At realtimeatwork.com/rtaw-pegase. provides a closed proprietary tool for the equivalent purpose of testing ECUs communications. Renault, for example, considers within their TSN tests, an ECU based on AUTOSAR classic software and interconnected with a physical and MAC Ethernet layer with TSN capable hardware [14]. The RTaW tool provides the Yet Another Next Generation (YANG) files for the hardware, with Bosch creating plugins to emulate detailed modeling of the ECUs integrate into the RTaW Pegase simulation [14].
Tools like Autoware [^83] At autowarefoundation.github.io. and Carla [^84] At carla.org. can then be used to simulate vehicular use cases, particularly in autonomous driving scenarios. Autoware in particular is considered a good solution both for being based on the widely used ROS and for offering hardware [^85] Available at autoware.org/autoware-open-ad-kit. kits that can be integrated within their software. However, while it has monitoring tools for CPU and GPU usage and some redundancy in sensor fault management, it lacks any in-vehicle networking, both wired and wireless [^86] Current focus of Autoware V2X module, details at tlab-wide.github.io/AutowareV2X..
CSS Electronics CANedge2878787Details at csselectronics.com. and equivalent equipment can be used to obtain real frames from CAN bus, either simulated or real, and send it to another computer. This can be used to obtain data to support the simulations with the aforementioned tools. If not possible, buses like CAN can also be simulated for example using Linux support for Virtual CAN tunnel (VXCAN). [^88] Source code available at github.com/torvalds/linux/blob/master/ drivers/net/can/vxcan.c..
7 Conclusions
In this article, the concept of Software Defined Vehicles was presented and framed as essential to support the development of the next generation of vehicular services. Considering the criticality of the future SDV services, as well as the increased complexity of development and deployment of services, it is more and more critical to consider a dynamic management of services.
The main building blocks of SDV to be considered in such a dynamic management were listed and detailed, as well as the main associated challenges. Vehicles will tend to a physical zonal architecture, supporting a set of services running in a set of higher performance ECUs (in a SOA paradigm). Existing operating systems will have to evolve towards interoperability and supporting virtualization and orchestration techniques, as well as integrating inter- and intra-vehicular communications with determinism techniques to support critical services that can have direct impact of the occupants and surroundings of the vehicles.
While the industry, including automakers and their suppliers, are working towards SDVs, their initiatives are still tending to be proprietary or closed but the trend is towards greater interoperability with holistic solutions such as ”the collaboration of collaborations” between AUTOSAR, COVESA, Eclipse SDV and SOAFEE announced in CES 2024.
A vision towards the dynamic management of services within SDV. Based on a deterministic network configurator, a data layer configurator, a orchestrator and a vehicle abstraction layer, all coordinated by a controller, it supports an easier dynamic management of vehicular services, easing the development burden of such artifacts.
Future work towards such vision includes the further specification, service orchestration, implementation and the proper evaluation of its scheduling algorithms, as well as an end-to-end acceptance tests.